### v1

# Webly-Supervised Image Manipulation Localization via Category-Aware Auto-Annotation

Abstract
[…]
Index Terms:
Image manipulation localization, automatic annotation, web supervision, image forensics.
I Introduction
The misuse of manipulated images can lead to fraud and the spread of misinformation, posing serious threats to social media security [1]. To this end, Image Manipulation Localization (IML) has gained increasing attention in recent years as a critical research problem [2, 3]. A central challenge in IML lies in the scarcity of large-scale high-quality data [4], which is essential for preventing model overfitting. Generating such data is extremely labor-intensive and time-consuming [5, 6]. It requires not only crafting convincing manipulated images but also providing precise pixel-level annotations of the forged regions. Synthetic data has been explored as an alternative [7, 8], but the resulting images often differ substantially from real-world forgeries, creating a significant domain gap that hinders model generalization [4]. Recent methods have attempted to close this gap by leveraging noise domain features [9, 10] and semantic suppression techniques [6], showing progress in improving robustness. Nevertheless, these methods do not fully resolve the fundamental issue: the lack of large-scale, manually forged data. As a result, overfitting remains a persistent challenge in IML.
Figure 1: When an image is tampered with by copy-pasting foreground objects from original images, the resulting original-tampered image pair belongs to Shared Donor Group (SDG). When an image is forged by directly modifying elements in place, the pair belongs to Shared Probe Group (SPG).
Figure 2: We propose a novel CIML method that processes the SPG and SDG images separately. It is used for automatic annotation, constructing a large-scale, high-quality dataset. The dataset notably enhances the generalization of image manipulation localization models. Additionally, the Object Jitter method is introduced to further enlarge the dataset size and data diversity.
Meanwhile, there are abundant manually forged images publicly available on the web, often accompanied by their corresponding originals. We propose to leverage such resources to address the data scarcity problem in IML. Specifically, we propose leveraging Constrained Image Manipulation Localization (CIML) models to automatically generate pixel-level mask annotations for these unlabeled forged images. Our key insight is that CIML simplifies the manipulation localization task by comparing a forged image with its corresponding original. This reduction in task complexity allows CIML to produce more accurate and reliable annotations for real-world manipulations. As a result, CIML offers a practical solution for automatically generating large-scale, high-quality training datasets for IML.
However, despite recent progress on less challenging data, existing CIML methods remain inadequate as reliable automatic annotators for complex real-world images, due to three major limitations. First, most prior approaches rely on a single correlation-based model to handle all input data [11, 12], which we argue is a suboptimal design. In practice, image pairs consisting of forged images and their corresponding original images can be categorized into two distinct types based on their relationship (see Fig. 1): (1) Shared Donor Group (SDG), where forged image is created by copying objects from the original image and pasting them onto another image. In SDG, the shared content between the real and fake images is the foreground object. (2) Shared Probe Group (SPG), where forged image is created by directly modifying the original image. In SPG, the shared content is the less object-specific background. Although correlation-based methods work reasonably well on SDG data, they are less effective on SPG data. This is because the shared backgrounds in SPG cover larger areas and offer fewer discriminative features compared to the localized, distinctive foregrounds in SDG. Training a single correlation model on both groups can therefore lead to confusion, resulting in degraded generalization and performance. Second, existing CIML methods overlook an important signal, that is the difference maps obtained by subtracting forged images from their originals. These difference maps consistently highlight manipulated regions and offer strong localization cues on SPG data. However, previous studies did not treat SPG and SDG images separately, thereby leaving this valuable signal underutilized. Third, CIML on SDG is inherently much more challenging than on SPG. Unlike SPG, CIML on SDG cannot benefit from image difference information. Additionally, SDG exhibits greater visual complexity and diverse manipulation patterns. Unfortunately, high-quality training data for SDG is extremely limited, making it difficult to train robust and generalizable models for real-world scenarios.
To tackle these problems, we propose a novel paradigm termed as Category-Aware Auto-Annotation v2 (CAAAv2), which treats image pairs in SDG and SPG separately. Our CAAAv2 comprises three components. At the beginning, a classifier is employed to determine whether an input image pair belongs to SDG or SPG. This classifier can be trained effectively through self-supervised learning using unlabeled images. Moving forward, a Difference-Aware Semantic Segmentation model that utilizes both the image pairs and their difference maps for accurate constrained manipulation localization in SPG. Additionally, a Correlation DINO model that improves generalization on SDG with a frozen DINOv2 [13] encoder, a Learnable Aggregation module, a Feature Super Resolution module and a Multi-Aspect Denoiser.
The proposed CAAAv2 is an upgraded version of our CAAA at CVPR 2024 [14]. In our earlier work, overfitting occurred in SDG because a trainable backbone was used during training with limited data. To address this issue, we now employ a frozen ViT backbone, which has resulted in significant improvements for CIML on SDG.
Subsequently, we collect a large amount of manually forged images from the Internet and then annotate their forged regions with the proposed CAAAv2. This approach considerably alleviates the data scarcity in image manipulation localization. To ensure the reliability of the annotations, we propose Quality Evaluation Score (QES), a novel metric that automatically evaluates annotation quality and filters out inadequate ones. Experiments demonstrate that our MIMLv2 dataset significantly improves generalization of various models on widely used benchmarks, as shown in Fig. 2.
Furthermore, the enhanced precision of CAAAv2 ensures that a greater number of high-quality samples are retained after filtering. Consequently, MIMLv2 is both larger and more diverse than our conference version MIMLv1 [14], better supporting the development of IML models.
To further enhance the scale and diversity of web supervision, we propose Object Jitter. This method generates semantically reasonable high-quality artifacts by introducing slight distortions to random objects, such as enlargement, overexposure, or texture alteration. Then, we train IML models to localize these distorted objects. Our approach effectively supplements MIMLv2 with additional authentic web images. This also differs from our conference paper that only uses MIMLv1 dataset to improve IML models.
To best leverage our high-quality web-scale supervision for IML, we further propose a new model named Web-IML. It comprises an encoder, a Multi-Scale Perception module, and a Self-Rectification module. The Multi-Scale Perception module enables the model to better integrate information from multiple perspectives. The Self-Rectification module allows the model to learn to identify and correct its mistakes. Our Web-IML significantly outperforms previous methods on various widely used benchmarks, being smaller yet more effective than the APSC-Net in our conference version.
In summary, our main contributions are as follows:
  • We propose to address the challenging IML task through the abundant web forgeries and their high-quality annotations distilled from a less challenging task, CIML.
  • We introduce a novel method CAAAv2 for CIML, along with a novel metric QES designed to assess prediction quality. Together, they are used to automatically annotate manually forged web images, constructing a large-scale, diverse, high-quality dataset MIMLv2. This dataset considerably alleviates the data scarcity in IML.
  • To further increase the size and diversity of the training data, we introduce Object Jitter, which generates high-quality artifacts using authentic web images.
  • We design a new model Web-IML, to better leverage the web-scale supervision. Extensive experiments have confirmed the effectiveness of the proposed method.
Figure 3: The overall pipeline of our proposed CAAAv2 paradigm. The input original-manipulated image pairs are first fed into a classifier to be classified into SPG or SDG. Then, we use Difference Aware Semantic Segmentation and Correlation DINO to predict forged regions in SPG and SDG respectively.
II Related Works
Image Manipulation Localization aims to identify forged regions within images [15, 16, 17, 18, 19] . Due to costly data collection, existing handcrafted datasets are tiny in size, leading to overfitting in many models [6, 4]. To address overfitting, some studies incorporated handcrafted features. RGB-N [9] utilized SRM [9] filters to suppress semantic bias. MVSS-Net [20] and MGQFormer [21] further exploited learnable noise domain modeling. CAT-Net [7], ObjectFormer [22] and HiFi-Net [23] leveraged frequency features to localize manipulations. However, the handcrafted features are mostly noisy and unstable, hindering further improvements in model performance. Recently, pure vision models with semantic suppression techniques have achieved better performance [24, 25]. NCL [4] employed contrastive learning to improve generalization. SparseViT [6] utilized dilated attention to reduce semantic bias. Despite the progress made, these works still suffer a lot from over-fitting due to scarce high-quality data. To this end, we propose to address data scarcity by constructing a large-scale high-quality dataset, with the Internet images accurately annotated in an automatic way.
Constrained Image Manipulation Localization aims at identifying the forged image region with reference to its original image [26]. Most prior works relied on correlation matching and treated SDG and SPG image pairs uniformly. DMVN [26] computed correlation maps to localize similar objects across image pairs. DMAC [11] incorporated dilated convolution for enhanced spatial information. AttentionDM [12] employed attention mechanism for better performance. MSTAF [27] performed correlation in both the encoder and decoder to extract better features. While these methods achieved notable progress on less challenging datasets (e.g., synthetic COCO [26]), their performance is limited in real-world scenarios characterized by high resolution, large variations, and complex details. In contrast to previous approaches, our method treats SPG and SDG separately, leveraging the differences between forged images and their originals. We are also the first to employ a frozen backbone for this task, effectively reducing overfitting.
III Category-Aware Auto-Annotation v2
For constrained image manipulation localization (CIML), previous works didn’t consider the critical discrepancy between SPG and SDG images, and processed them using a single correlation-based model. We argue that such a paradigm is sub-optimal and the reasons are as follows:
First, the similar regions in SDG images are typically foreground objects (e.g., the cats in the SDG branch of Fig. 3) that have specific, similar shapes and distinctive features. In contrast, the similar regions in SPG images are background, which generally lacks distinctive features for accurate correlation matching (e.g., for the images of the SPG branch in Fig. 3, a patch of snow in the background is similar to all other snow patches). These background regions will cause confusion for correlation-based models, especially in complex scenarios.
Second, the difference between the forged image and its original in an SPG pair serves as an important cue. Most areas in an SPG image pair are almost identical and spatially aligned (e.g., the image pair in the SPG branch of Fig. 3). Simply subtracting between them yields a difference map that can always highlight the manipulated regions. However, this information is difficult to utilize in correlation-matching-based models and was therefore overlooked in prior works.
Based on these observations, we propose a novel paradigm for the CIML task, Category-Aware Auto-Annotation v2 (CAAAv2). The key idea is to process SPG and SDG images independently, as shown in Fig. 3. First, a classifier (described in Section III.A) categorizes the input image pairs as either SPG or SDG. For SPG pairs, we apply Difference-Aware Semantic Segmentation (Section III.B). For SDG pairs, we process the images using Correlation DINO (Section III.C). Importantly, the trained models are further used to auto-annotate unlabeled forged images, thereby addressing the data scarcity issue in IML tasks.
The difference between CAAAv2 and its conference version CAAA lies in the SDG branch model. In CAAA, we processed SDG image pairs using a trainable CNN structure, which tended to overfit the limited training data. In contrast, CAAAv2 employs a well pre-trained frozen ViT backbone, thereby reducing overfitting to task-specific data. Moreover, the ViT backbone naturally integrates global information, thereby alleviating semantic misalignment caused by scale diversity.
III-A Self-Supervised Classification
To achieve the classification of SPG and SDG, we propose training a classifier via self-supervised learning on unlabeled images. Given an image, we apply random augmentations (e.g. color jitter, image compression) and manipulations (copy-paste, splicing or removal). The manipulated image and the original image form an SPG pair. To construct an SDG pair, we copy 1-3 random regions from the original image, resize them, and paste them into another image. With the obtained image pairs, we can effectively train our classifier. Our classifier actually only needs to discern whether the main area of the input image pair is identical or not, without considering which image or region is fake. For example, in Fig. 8, the two images in any SPG pair are nearly identical and spatially aligned, whereas images in any SDG pair are distinctly different. These characteristics are universally applicable and strong enough. Therefore, this classification task is quite simple, enabling accurate separation of images into the two groups.
Figure 4: Manipulated images often degrade during transmission. As a result, the absolute difference between a tampered image and its original does not accurately indicate the forged region. Our method effectively denoises the difference map by leveraging semantic information.
III-B Difference-Aware Semantic Segmentation (DASS)
Ideally, for an SPG image pair, the absolute difference between the authentic and forged images would directly reveal the forged region. However, due to transmission noise and image degradation [28], the absolute difference map is often imprecise. As shown in Fig. 4, almost every area in the difference map shows nonzero values because of noise. Even after binarizing the difference map using an OTSU [29] adaptive threshold, noisy highlights persist, especially in high-frequency regions such as edges. To address this issue, we propose denoising the difference map by incorporating semantic information from the images. Specifically, we concatenate the authentic image, the forged image, and their absolute difference map along the channel dimension, and feed the result into a semantic segmentation model. Our segmentation model employs a VAN [30] encoder for semantic feature extraction and utilizes the same Multi-Aspect Denoiser as in our Corr-DINO to denoise the difference map.
III-C Correlation DINO (Corr-DINO)
CIML on SDG requires detecting shared objects at various scales within complex backgrounds. This task is considerably more challenging than SPG, where targets are already highlighted through image difference maps. As a result, synthetic or small-scale handcrafted datasets are largely insufficient for training a robust model for complex real-world scenarios.
To tackle the overfitting issue, we introduce a novel frozen-denoising paradigm (see Fig. 5). Our key insight is that overfitting can be mitigated by employing a frozen encoder and subsequently denoising its output features. Building on this paradigm, we developed the Corr-DINO model (see Fig. 6), which consists of five components:
1. Frozen Encoder for feature extraction. We extract features from both the tampered image $I_{a}$ and its original image $I_{b}$ with a frozen DINOv2 [13] backbone. This ViT backbone has learned generalized representations through self-supervised learning on vast-scale web images. The features from the last ViT layer, denoted as $F_{a1}$ and $F_{b1}$, contain rich semantic information and are used to compute correlation features.
2. Correlation module for correlation features calculation. $F_{a\_corr}=[Corr(F_{a1},F_{b1}),Corr(F_{a1},F_{a1})]$, $F_{b\_corr}=[Corr(F_{b1},F_{a1}),Corr(F_{b1},F_{b1})]$. The same correlation function $Corr$ as most previous works [12, 27, 14] is adopted. The detailed $Corr$ function is presented in the Appendix.
Figure 5: Different from previous works that employ learnable encoders and rule-based feature aggregation for constrained image manipulation localization on SDG image pairs, we propose a novel frozen-denoising paradigm. This paradigm offers significantly better generalization in real-world scenarios.
Figure 6: The overall pipeline of our proposed Correlation DINO. First, we employ a frozen DINO backbone to extract image features. The features obtained from the final DINO layer are then used for correlation calculation. Next, a learnable aggregation module reduces the channel dimension of the correlation features, which are further refined using a Feature Super Resolution module. Finally, the refined features are denoised with a Multi-Aspect Denoiser, resulting in the final mask prediction. Green lines indicate operations used only during training with synthetic data, due to the lack of mask labels for the real images.
3. A Learnable Aggregation module for dimension reduction. Previous works simply select the top-K activated channels with the highest average values in $F_{a\_corr}$ and $F_{b\_corr}$ to reduce dimension. This approach works for trainable encoders but is suboptimal for our frozen encoder, whose correlated object features are relatively sparse. Selecting only the top-K channels may discard important features and compromise the integrity of $F_{a\_corr}$ and $F_{b\_corr}$. To overcome this, we propose a learnable aggregation module that uses trainable parameters to reduce the channel count to K. Following prior methods, we also include statistics such as average and maximum values across the channel dimension. $F_{a\_aggr}=[C_{1\times 1}(R(C_{1\times 1}(F_{a\_corr}))),Avg(F_{a\_corr}),Max(F_{a\_corr})]$. Here, $C_{1\times 1}$ is $1\times 1$ conv-layer and $R$ is ReLU. A similar process is used for $F_{b\_corr}$ to generate $F_{b\_aggr}$.
4. Feature Super Resolution module for image detail reconstruction. The features from the last ViT layer are rich in semantics but lack fine image details [13], making $F_{a\_aggr}$ and $F_{b\_aggr}$ too coarse for fine-grained mask predictions. To reconstruct image details, we leverage features from the last 2-4 layers of DINO, which are known to be more sensitive to edges and fine details [13]. We concatenate the features from the last four encoder layers $F_{a}=Concat(F_{a1,2,3,4})$ and interpolate $F_{a\_aggr}$ to different scales ($1\times,2\times,4\times,8\times$) to obtain $F_{aggr\_a1,2,3,4}$. We then fuse them with $1\times 1$ conv-layers. $F_{asr\_i}=Conv_{1\times 1}(Conv_{1\times 1}(F_{a})+Conv_{1\times 1}(F_{aggr\_ai}))$ for iin(1,2,3,4)i\,in\,(1,2,3,4). A similar process is applied to $F_{b\_aggr}$.
5. Multi-Aspect Denoiser module for features denoising. Since the ViT encoder is frozen, its output features inevitably contain significant task-irrelevant noise. To achieve precise mask prediction, we propose denoising the correlation features by integrating multi-perspective information. This module fuses multi-scale features $F_{asr\_1,2,3,4}$ in a top-down manner. $F^{\prime}4=Conv_{3\times 3}([F_{asr\_4},Avg(F_{asr\_4})])$, $F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{asr\_n}))$ for iin(1,2,3)i\,in\,(1,2,3). Here, $Conv_{3\times 3}$ and $Conv_{1\times 1}$ are conv-layers with kernel size 3 and 1, and $Avg$ is global average pooling. The fused features are further concatenated and denoised using a set of dilated $3\times 3$} conv-layers. The output is concatenated to $F^{\prime}1$ and processed by a $3\times 3$} conv-layer to yield the mask prediction.
Our Correlation DINO is optimized end-to-end using cross-entropy loss. It significantly alleviates the overfitting problem in challenging real-world CIML tasks for SDG.
IV MIMLv2 Dataset
In this section, we introduce MIMLv2, a large-scale, diverse, and high-quality dataset. The key idea is to leverage constrained image manipulation localization models trained on existing datasets to automatically annotate manually forged images collected from the web. To ensure high annotation quality, we propose a novel metric that filters out inadequate annotations.
Our MIMLv2 is annotated using our CAAAv2, which is significantly more accurate than its conference version, CAAA [14]. As a result, more samples are retained as high quality after filtering, as shown in Table I. Compared to MIMLv1 [14], MIMLv2 offers greater scale and diversity, making it a more effective resource for training IML models.
Figure 7: The overall construction pipeline of our MIMLv2 dataset.
IV-A Dataset Construction
As illustrated in Fig. 7, we construct MIMLv2 through the following steps:
Image Collection. We collect forged images and their corresponding originals from imgur.com. On this website, the images are manually manipulated and thus have high-quality, diverse forged regions. Additionally, we exclude images that already appear in the evaluation datasets (Section VII) and remove duplicates by checking their MD5 and pHash.
Auto-Annotation. We utilize the CAAAv2 models, as proposed in Section III, to obtain mask annotations for the collected images. These models are sufficiently trained with one million synthetic COCO samples [26], along with the handcrafted datasets CASIAv2 [31] and IMD20 [32].
[…]
Figure 8: Some images and their corresponding mask annotations from the proposed MIMLv2 dataset. Left: SPG image pairs. Right: SDG image pairs.
IV-B Dataset Highlights
We present a few samples of the proposed MIMLv2 dataset in Fig. 8. The main highlights of our dataset are as follows:
  • High quality. Image manipulations in the proposed dataset are meticulously crafted by humans. Such data can teach models to detect real-world forgery, rather than merely overfitting to simple patterns found in synthetic data.
  • Large Scale. As shown in Table I, the proposed dataset comprises 246,212 manually forged images. This is dozens of times more samples than previous handcrafted IML datasets (e.g., $\approx$120 times more samples than IMD20 [32]).
  • Board Diversity. MIMLv2 comprises images of various sizes, styles, and manipulation types (e.g., copy-move, splicing, removal, AIGC). These images are created by hundreds of thousands of individuals using various software. Such broad diversity can considerably enhance the generalization ability of deep IML models.
  • Modern Style. MIMLv2 contains abundant modern images that were recently captured and forged, reflecting current photography and forgery technologies. In contrast, the CASIA dataset [31] proposed over a decade ago, features blurred small images and outdated forgeries. Therefore, MIMLv2 better meets the current requirements for IML.
  • Strong Scalability. With the rising popularity of online image manipulation competitions (e.g., PS-Battles [33], which attracts over 20 million participants), new manually forged images are continuously generated. Our dataset construction approach is ready to harness this growing cheap web data. Therefore, our dataset can be easily expanded in the future.
TABLE I: Comparison between different handcrafted image manipulation localization datasets. ’Num.’ denotes number.
Name Year Real Num. Forged Num. (Height, Width) Range
CASIAv1 [31] 2013 800 921 (246, 384)-(500, 334)
CASIAv2 [31] 2013 7461 5,123 (160, 240)-(901, 600)
Coverage [34] 2016 100 100 (190, 334)-(472, 752)
NIST16 [35] 2016 224 564 (500, 500)-(3744, 5616)
In Wild [36] 2018 0 201 (650, 650)-(2736, 3648)
IMD20 [32] 2020 414 2,010 (193, 260)-(4437, 2958)
MIMLv1 [14] 2024 0 123,150 (45, 120)-(13846, 9200)
MIMLv2 (Ours) 2025 63,847 246,212 (45, 120)-(13846, 9200)
V Object Jitter
We experimentally find that the IML model performance can be further improved with more high-quality data. To this end, we introduce Object Jitter, an effective method that enhances IML models by generating high-quality image artifacts online during training from authentic web images.
Previous works synthesize tampered images by randomly cropping and pasting image objects [37, 9]. However, these data synthesis methods often produce overly obvious artifacts and semantically unreasonable content, as shown on the left in Fig. 9. Models trained on such data typically overfit to simple cases and perform poorly on real-world forgeries, where artifacts are subtle and tampered content is semantically reasonable. DEFACTO [8] attempts to ensure semantic reasonableness by designing elaborate rules. It manually identifies eight object types with consistently similar shapes and appearances (e.g., the road sign in the middle of Fig. 9); copy-move or splicing operations are restricted to objects of the same type. Because DEFACTO applies only to a limited number of object classes, the generated data lacks diversity, causing models to overfit the semantics of these specific object types. Furthermore, DEFACTO heavily relies on accurate object class annotations, which significantly limits its broader application.
To overcome these challenges, we propose a novel idea: instead of altering image semantics through tampering, we simulate subtle artifacts. Building on this idea, we introduce Object Jitter, a novel method detailed in Fig 10. Given an input image, we readily obtain a segmentation mask for a random object using off-the-shelf segmentation models like SAM [38]. The selected object is then processed using one or a combination of the following three random operations:
  • Size Jitter. Slightly enlarge the selected object’s region and its corresponding mask by a random scale.
  • Exposure Jitter. Slightly overexpose the selected object.
  • Texture Jitter. Slightly alter the texture of the selected object region by applying a random combination of JPEG compression, reverse JPEG compression [39], and blur.
After processing, we add a random smooth transition on the edges to minimize visual artifacts. We apply Object Jitter to 1 to 3 random objects within an input image. The mask regions of these processed objects are recorded and subsequently considered as tampered regions during training.
Our method successfully addresses existing issues: First, it always preserves the semantics of the input image, thereby preventing models from overfitting to simple cases with abnormal semantics. Second, by applying subtle changes to objects that already naturally blend into the background, the generated visual image artifacts are unobvious. Third, our method is unconstrained and universally applicable. Consequently, Object Jitter consistently produces high-quality artifacts and effectively supplements our MIMLv2 dataset.
Figure 9: Comparison between different data generation methods. Our method is universally applicable and always produces semantically reasonable forgeries.
Figure 10: The proposed Object Jitter method. Given a randomly selected object from an authentic image, we slightly enlarge or overexpose it, or alter its texture. The processed object is regarded as tampered object for training.
VI Web-IML
To better leverage the web-scale supervision, we propose a new model called Web-IML. As illustrated in Fig. 11, Web-IML comprises a feature extractor, a Multi-Scale Perception module to integrate information from multiple perspectives, and a Self-Rectification module that enables the model to identify and correct its mistakes.
Figure 11: The overall framework of the proposed Web-IML model.
Multi-Scale Perception module motivation. During meticulous image forensic analysis, humans often zoom in and out the image repeatedly, comparing between diverse observations to assist their final prediction. To mimic the human perception way, we introduce the Multi-Scale Perception module.
Multi-Scale Perception module pipeline. Specifically, given four feature maps extracted from the backbone model $F_{1},F_{2},F_{3},F_{4}$}, we first extract global features in different scales: $G_{i}=Conv_{1\times 1}(Avg_{i}(F_{4}))\,for\,i\,in\,(1,2,3,6)$, where $Conv_{1\times 1}$ denotes conv-layer with kernel size 1 and $Avg_{i}$ denotes average pooling function with output size $i\times i$. The global features are then resized, concatenated to $F_{4}$ and fused, dimension reduced to get $F^{\prime}4$. $F^{\prime}4=Conv_{3\times 3}([G_{1},G_{2},G_{3},G_{6},F_{4}])$. Next, the high-level features are fused with the low-level features recursively. $F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{n}))\,for\,i\,in\,(1,2,3)$. This process fully integrates high-level semantics with low-level image details, resulting in the multi-scale features $F^{\prime}_{1},F^{\prime}_{2},F^{\prime}_{3},F^{\prime}_{4}$}.
Self-Rectification module motivation. Due to the challenging nature of the image manipulation localization task, forensic models inevitably overlook key information and make mistakes. To minimize negligence, we propose the Self-Rectification module. With this module, our model can learn to check and correct its initial predictions.
Self-Rectification module pipeline. As shown at the bottom of Fig. 11, we obtain the initial mask prediction $P_{n}$ from $F^{\prime}_{1}$ via a $\times1$ conv-layer. We then adaptively obtain a rectification feature $F_{r}$ through a tiny Rectification Feature Mapping module ($RFM$), which consists of 5 cascaded conv-layers. $F_{r}=RFM(P_{n})$. The $F_{r}$ informs the model of suspected tampered regions and potentially incorrect prediction regions. Then $F_{r}$ is fused with $F_{a}$ with a $\times3$ conv-layer to get $F_{c}$. $F_{c}=Conv_{3\times 3}([F_{a},P_{c}])$. To enable in-depth, meticulously analyses of $F_{c}$, we introduce the Nested Channel Attention module. As shown in the top left of Fig. 11, it can be formulated as $G_{1}=Avg(F_{c})W_{1}$, $G_{2}=G_{1}W_{2}$, $A_{2}=\sigma(W_{3}G_{2})$, $A_{1}=\sigma(W_{4}G_{1}*A_{2})$, $F_{o}=F_{c}*A_{1}$. Here, $G_{1,2}$ are global vectors, $A_{1,2}$ are attention maps, $W_{1,2,3,4}$ are learnable parameters, $Avg$ is global average pooling and $\sigma$ is the sigmoid function. The output $F_{o}$ is further processed by four $\times3$ conv-layers with dilation sizes $1,2,3,6$ respectively, the output features are concatenated and fused with a $\times3$ conv-layers. The mask prediction $P_{f}$ is obtained from $F_{o}$ with a $\times1$ conv-layer. To get the refined final prediction $P_{f}$, we replace $P_{n}$ with $P_{f}$, replace $F_{a}$ with $F_{o}$, then repeat the self rectification once again.
Our Web-IML is optimized with the cross entropy loss calculated between the prediction mask and the ground-truth.
TABLE II: Comparison on image manipulation localization. ”WS” is our proposed web supervision, which trains models using data from the MIMLv2 and Object Jitter. ”Avg*” is the average score excluding IMD20. ”Param.” is the number of model parameters.
Method CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] CocoGlide [10] CIMD [40] MISD [41] Avg.* Average Param.
IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
ManTra-Net [42] .086 .130 .181 .271 .040 .062 .098 .146 .155 .203 .058 .096 .105 .171 .104 .156 .103 .154 8M
RRU-Net [43] .330 .380 .165 .260 .080 .129 .169 .256 .223 .304 .036 .068 .440 .570 .212 .285 .206 .281 5M
MVSS-Net [20] .403 .435 .389 .454 .243 .294 .200 .260 .276 .357 .016 .021 .120 .175 .235 .284 .235 .285 146M
PSCC-Net [37] .410 .463 .340 .446 .067 .110 .115 .192 .333 .422 .212 .290 .400 .491 .294 .370 .268 .345 4M
CAT-Net [7] .684 .738 .238 .292 .238 .302 - - .290 .366 .282 .344 .314 .394 .341 .406 - - 114M
IF-OSN [28] .465 .509 .181 .268 .247 .326 .259 .364 .207 .264 .195 .295 .375 .464 .278 .354 .341 .406 124M
EVP [44] .438 .502 .078 .114 .188 .239 .177 .268 .084 .118 .224 .283 .392 .531 .234 .298 .276 .356 63M
TruFor [10] .630 .692 .446 .522 .279 .348 - - .294 .362 .322 .418 .426 .549 .400 .482 - - 67M
PIM [25] .512 .566 .188 .251 .225 .280 .340 .419 .327 .404 .177 .229 .389 .519 .303 .375 .308 .381 168M
APSC-Net [14] (Conf.) .799 .837 .490 .523 .398 .436 .339 .391 .243 .283 .517 .565 .420 .524 .477 .528 .458 .508 143M
Web-IML (Ours) .823 .857 .508 .551 .412 .465 .370 .422 .285 .339 .629 .673 .430 .545 .515 .572 .494 .550 140M
APSC-Net (w/ MIMLv1) .810 .848 .498 .568 .525 .590 .679 .760 .292 .355 .652 .720 .484 .587 .543 .611 .562 .632 143M
Web-IML (w/ MIMLv1) .821 .854 .561 .616 .520 .587 .692 .768 .275 .325 .789 .845 .496 .610 .578 .640 .594 .658 140M
APSC-Net (w/ WS) .832 .865 .594 .642 .578 .661 .690 .766 .389 .451 .660 .725 .485 .588 .598 .655 .604 .671 143M
Web-IML (w/ WS) .846 .879 .715 .768 .591 .670 .698 .776 .439 .507 .759 .821 .495 .612 .641 .709 .649 .719 140M
TABLE III: Ablation study on our Web-IML model. ‘MP‘ is the Multi-Scale Perception module. ‘SR‘ is the Self-Rectification module. ‘NA‘ is our Nested Channel Attention. ‘RF‘ denotes the second round self-rectification. ‘WS‘ denotes the proposed Web Supervision.
Num. Ablation Settings CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] CocoGlide [10] CIMD [40] MISD [41] Average
MP SR NA RF WS IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
(1) $\times$ $\times$ $\times$ $\times$ $\times$ .711 .779 .361 .430 .346 .410 .273 .342 .345 .417 .262 .338 .362 .471 .380 .455
(2) $\times$ $\times$ $\times$ $\times$ .775 .814 .440 .487 .359 .415 .351 .406 .275 .324 .450 .502 .375 .482 .432 .490
(3) $\times$ $\times$ $\times$ .802 .836 .523 .576 .385 .441 .364 .415 .287 .343 .557 .610 .382 .493 .471 .530
(4) $\times$ $\times$ .816 .848 .549 .598 .394 .449 .369 .419 .291 .345 .593 .649 .386 .496 .486 .543
(5) $\times$ .823 .857 .508 .551 .412 .465 .370 .422 .285 .339 .629 .673 .430 .545 .494 .550
(6) .846 .879 .715 .768 .591 .670 .698 .776 .439 .507 .759 .821 .496 .612 .649 .719
TABLE IV: Robustness evaluation with AUC metric on NIST16 dataset.
Resize Blur JPEG
Method Ori .78x .25x k=3 k=15 q=100 q=50
ManTra-Net [42] .795 .774 .755 .774 .746 .779 .744
SPAN [45] .840 .832 .802 .831 .792 .836 .807
PSCC-Net [37] .855 .853 .850 .854 .800 .854 .854
ObjectFormer [22] .872 .872 .863 .860 .803 .864 .862
Sparse ViT [6] .888 .884 .869 .881 .877 .886 .881
ERMPC [46] .895 .893 .877 .892 .871 .894 .888
UnionFormer [47] .881 .873 .872 .865 .843 .880 .880
APSC-Net+MIMLv1 .928 .917 .888 .907 .900 .922 .907
Web-IML (Ours) .942 .935 .904 .914 .900 .940 .909
TABLE V: Ablation study on our Web Supervision method with Web-IML model. ‘PSCC-Synthetic‘ denotes the synthetic data from the PSCC-Net [37]. ‘O.J.‘ denotes the proposed Object Jitter method. ‘C.G.‘ denotes the CocoGlide [10]. $\Delta$‘ denotes improvement.
CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] C.G. [10] CIMD [40] MISD [41] Average ($\Delta$)
Num. Training Data IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
(1) Baseline .823 .857 .508 .551 .412 .465 .370 .422 .285 .339 .629 .673 .430 .545 .494 (+0.0%) .550 (+0.0%)
(2) +PSCC-Synthetic [37] .813 .845 .503 .546 .348 .392 .361 .410 .177 .211 .598 .654 .330 .441 .447 (-9.5%) .500 (-9.1%)
(3) +DEFACTO [8] .809 .843 .354 .380 .412 .465 .365 .424 .359 .418 .394 .438 .461 .572 .451 (-8.7%) .506 (-8.0%)
(4) +AIGC-Synthetic [48] .817 .849 .273 .296 .220 .276 .284 .337 .394 .462 .287 .340 .373 .429 .376 (-24.%) .420 (-23.%)
(5) +MIMLv1 [14] (Conf.) .821 .854 .561 .616 .520 .587 .692 .768 .275 .325 .789 .845 .496 .610 .594 (+20.%) .658 (+20.%)
(6) +MIMLv2 .849 .883 .694 .755 .518 .584 .697 .772 .376 .435 .716 .776 .448 .564 .614 (+24.%) .681 (+24.%)
(7) +MIMLv1 [14]+O.J. .833 .867 .587 .643 .549 .620 .693 .768 .344 .407 .793 .849 .487 .602 .612 (+24.%) .679 (+23.%)
(8) +O.J. .822 .853 .575 .616 .270 .310 .422 .482 .308 .362 .698 .760 .413 .521 .501 (+1.4%) .558 (+1.4%)
(9) +MIMLv2+O.J. (Ours) .846 .879 .715 .768 .591 .670 .698 .776 .439 .507 .759 .821 .496 .612 .649 (+31.%) .719 (+31.%)
TABLE VI: Ablation study on our Object Jitter method. ‘COCO‘ denotes using the COCO [49] dataset images for Object Jitter. ’SA’ denotes using 400k random images from the SA-1B [38] dataset for Object Jitter. ‘C.G.‘ denotes the CocoGlide dataset [10].
Num Object Jitter Operation
Image
Source
CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] C.G. [10] CIMD [40] MIFD [41] Average
Size Texture Exposure IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
(1) $\times$ $\times$ $\times$ SA .849 .883 .694 .755 .518 .584 .697 .772 .376 .435 .716 .776 .448 .564 .614 .681
(2) $\times$ SA .839 .872 .634 .687 .528 .592 .703 .778 .396 .456 .748 .798 .487 .605 .619 .684
(3) $\times$ SA .844 .878 .669 .729 .554 .618 .701 .778 .427 .491 .780 .834 .501 .616 .640 .706
(4) $\times$ SA .839 .872 .699 .754 .560 .627 .699 .774 .413 .479 .802 .861 .512 .628 .646 .714
(5) $\times$ $\times$ SA .841 .875 .673 .732 .540 .602 .701 .777 .432 .496 .758 .816 .516 .631 .637 .704
(6) COCO .848 .881 .645 .711 .584 .657 .699 .776 .489 .549 .803 .864 .496 .612 .652 .721
(7) SA .846 .879 .715 .768 .591 .670 .698 .776 .439 .507 .759 .821 .496 .613 .649 .719
VII Experiments
VII-A Experiments on the Web-IML Model (IML Task)
This subsection evaluates the effectiveness of the proposed Web-IML model by comparing it with other IML models on multiple real-world handcrafted and AIGC forgery benchmarks. We also perform an ablation study to provide an in-depth analysis of the Web-IML model.
Implementation Details. For Web-IML, we adopt ConvNeXt-Base [50] as the feature extractor. The model is trained for 200k iterations with a batch size of 16, using a $\times512$ input size, consistent with previous works [7, 10]. We employ Cross-Entropy loss and AdamW [14] optimizer, with the learning rate linearly decaying from 1e-4 to 1e-6. CASIAv2 [31] and tampCOCO [7] serve as baseline training sets. For fair comparison, mask predictions are binarized with a fixed threshold of 0.5 and evaluated using IoU and binary F1-score.
Comparison Study. Table II shows that our Web-IML significantly outperforms existing methods on multiple real-world image forgery benchmarks. For example, it achieves an average IoU that is 24.1 points higher than TruFor. Furthermore, Web-IML consistently outperforms our conference version, APSC-Net, across various settings. These results demonstrate the strong generalization capabilities of our Web-IML model. Qualitative comparison is illustrated in Fig. 12.
Ablation Study. The ablation results are shown in Table III. Setting (2), which includes the proposed Multi-Scale Perception (MP) module, achieves 5.2 points higher IoU than the baseline (1). This indicates that the proposed MP module effectively integrates high-level and low-level features to detect diverse image artifacts. Setting (3), which further incorporates the Self-Rectification (SR) module, yields an additional 3.9 point IoU improvement over Setting (2), as SR enables the model to detect and correct its errors. Setting (5) outperforms (4) via our nested attention module that enables in-depth analysis. Setting (6) outperforms (5) by including a second round of self-rectification refinement, demonstrating that further refinement can minimize incorrect predictions.
Robustness Evaluation. We evaluate Web-IML under various distortions, including image resizing (scale factors 0.78 and 0.25), Gaussian blur (kernel sizes 3 and 15), and JPEG compression (quality factors 50 and 100). The results are shown in Table IV. Our model exhibits stable performance under these conditions, demonstrating strong robustness.
VII-B Experiments on the Web-Supervision Method (IML Task)
This subsection evaluates the effectiveness of our proposed web-supervision approach in addressing the data scarcity issue for IML. Our web-supervision encompasses both the MIMLv2 dataset and the Object Jitter method. We compare the performance of the same IML model trained on data created by our web-supervision approach against other synthetic methods.
Implementation Details. The implementation details remain the same as in the previous subsection VII.A. When supplementing the baseline training sets with additional IML training sets, the total training volume is kept constant with a uniform sampling ratio across all datasets.
Comparison Study for the MIMLv2 Dataset. Table V presents the IML performance of the Web-IML model when trained on different datasets. The baseline (1) uses the two commonest training sets: tampCOCO [7] and CASIAv2 [31]. Incorporating synthetic training data from PSCC-Net [37], DEFACTO [8] or OpenSDID [48] diminishes the average model performance (settings 2, 3, 4). This is because the first two synthetic datasets do not notably surpass the baseline data in quality or diversity, while the AIGC forgeries from OpenSDID differ significantly from non-AIGC forgeries. These results confirm that merely increasing training data scale does not lead to improved performance. In contrast, including the MIMLv1 (setting 5) or MIMLv2 (setting 6) datasets significantly improves model performance by an average of 10 and 12 points, respectively, without increasing training or inference burden. These results confirm that our MIMLv1 and MIMLv2 datasets considerably alleviate data scarcity and overfitting issues in real-world scenarios. This is attributable to their large scale, manual forgery, and considerably higher quality and diversity compared to other datasets. Notably, MIMLv2 yields greater improvement due to its larger quantity of high-quality samples and superior diversity compared to MIMLv1.
Comparison Study for the Object Jitter. Table V shows that including the training data generated by our Object Jitter method further improves the model when combined with MIMLv1 (setting 7 outperforms 5) or MIMLv2 (setting 9 outperforms 6) datasets, effectively supplementing them. This is attributed to our Object Jitter method’s consistent generation of diverse, high-quality samples, which further enriches the training data.
Ablation Study for the Object Jitter. Table VI presents the ablation study for Object Jitter operations and image source. Setting (1) is the baseline, where Web-IML is trained with MIMLv2 but without Object Jitter. Setting (7) represents the full Object Jitter method. Settings (2, 3, 4) involve removing the size jitter, texture jitter, and exposure jitter operations, respectively; all of them result in worse performance than setting (7). These results demonstrate that all three operations contribute to improved performance through increased data diversity. Notably, Setting (2), which lacks size jitter, is significantly worse than (7). Meanwhile, Setting (5), which uses only size jitter, performs close to (7). This indicates that the size jitter operation is the primary contributor to the improvement. Since most Internet forgeries are not generated by copy-move, they exhibit exposure and texture artifacts. The size jitter operation, without altering the object’s exposure or texture, reduces model’s reliance on exposure or texture cues, thus better supplementing MIMLv2. Furthermore, Setting (6) applies the full Object Jitter method using COCO [49] as the image source, instead of 400k random images from SA [38] in setting (7). The similar performance of Settings (6) and (7) demonstrates that the Object Jitter method is robust to the choice of image source.
Conclusions Verification on Other IML Models. We extended our verification process to other IML models. Specifically, we evaluated MIMLv2 and Object Jitter on PSCC-Net [37], CAT-Net [7], and TruFor [10] using their official model code. Table VII presents the results, which clearly indicate that MIMLv1, MIMLv2, and Object Jitter all lead to significantly higher scores. MIMLv2, in particular, consistently yields significantly greater improvement than MIMLv1 (conference version). Moreover, Object Jitter further boosts performance when integrated with MIMLv2. These results robustly confirm our previous analyses and conclusions.
TABLE VII: Comparison study of our web-supervision on IML task. Blue number for improvement and Red number for decrease. ’B.L.’ denotes baseline setting using tampCOCO [7] and CASIAv2 [31] datasets as training data. ’PSCC-Synth.’ denotes the synthetic dataset of PSCC-Net [37], ’AIGC-Synth.’ denotes the AIGC forgery dataset OpenSDID [48]. ’O.J.’ denotes our Object Jitter method.
Training Data CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] CIMD [40] Average
IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$
PSCC-Net B.L. .401 0 .430 0 .197 0 .218 0 .247 0 .295 0 .125 0 .156 0 .161 0 .221 0 .226 0 .264 0
+PSCC-Synth. .352 12% .379 12% .147 25% .159 27% .230 7% .274 7% .101 19% .125 20% .101 37% .124 44% .186 18% .212 20%
+DEFACTO[8] .394 2% .429 1% .231 17% .256 17% .223 10% .270 8% .157 26% .171 10% .120 25% .159 28% .225 0% .257 3%
+AIGC-Synth. .346 16% .377 14% .109 81% .123 77% .091 171% .142 108% .110 14% .144 8% .124 30% .170 30% .156 45% .191 38%
+MIMLv1 (Conf.) .609 52% .649 51% .395 101% .477 119% .402 63% .476 61% .470 276% .541 247% .206 28% .251 14% .416 84% .479 81%
+MIMLv2 .602 50% .642 49% .402 104% .498 128% .405 64% .485 64% .575 360% .662 324% .354 120% .483 119% .468 +107% .554 110%
+MIMLv2+O.J .609 52% .651 51% .451 129% .550 152% .406 64% .486 65% .570 356% .658 322% .335 108% .464 110% .474 110% .562 113%
CAT-Net B.L. .660 0 .703 0 .245 0 .286 0 .239 0 .287 0 .157 0 .192 0 .229 0 .263 0 .306 0 .346 0
+PSCC-Synth. .410 38% .437 38% .215 12% .232 19% .288 21% .331 15% .121 23% .148 23% .160 30% .189 28% .239 22% .268 23%
+DEFACTO[8] .673 2% .715 2% .200 18% .230 20% .220 8% .261 9% .164 4% .200 4% .130 43% .156 41% .277 9% .312 10%
+AIGC-Synth. .646 2% .685 3% .136 80% .147 95% .170 41% .242 19% .130 21% .165 16% .088 160% .117 125% .234 31% .271 28%
+MIMLv1 (Conf.) .691 5% .728 4% .302 23% .389 36% .353 48% .422 47% .547 248% .629 228% .397 73% .477 81% .458 50% .529 53%
+MIMLv2 .684 4% .720 2% .368 50% .443 55% .338 41% .416 45% .562 258% .644 235% .470 105% .548 108% .484 58% .554 60%
+MIMLv2+O.J .703 7% .742 6% .355 45% .437 53% .358 50% .440 53% .575 266% .656 242% .497 117% .575 119% .498 63% .570 65%
TruFor B.L. .755 0 .794 0 .363 0 .411 0 .320 0 .377 0 .248 0 .289 0 .296 0 .346 0 .396 0 .443 0
+PSCC-Synth. .750 1% .798 1% .340 7% .384 7% .325 2% .373 1% .253 2% .298 3% .278 6% .331 5% .389 2% .437 2%
+DEFACTO[8] .752 0% .791 0% .322 13% .357 15% .214 50% .260 45% .209 19% .247 17% .184 61% .217 59% .336 18% .374 18%
+AIGC-Synth. .746 1% .791 0% .220 65% .261 57% .213 50% .263 43% .221 12% .267 8% .121 146% .156 122% .304 30% .348 28%
+MIMLv1 (Conf.) .755 0% .794 0% .385 6% .458 10% .352 9% .408 8% .596 58% .678 57% .340 13% .401 14% .486 18% .548 19%
+MIMLv2 .767 2% .794 0% .387 6% .458 10% .407 21% .472 20% .618 60% .696 58% .370 20% .422 18% .510 22% .568 22%
+MIMLv2+O.J .770 2% .811 2% .466 22% .511 20% .439 27% .502 25% .619 60% .700 59% .309 4% .357 3% .521 24% .576 23%
TABLE VIII: Comparison study of CIML on SPG (left) and SDG (right). ’*’ denotes model trained with both SDG and SPG data. ’Nonzero’ denotes using the non-zero region of the difference map as mask, ’OTSU’ denotes the difference map binarized by OTSU. ’QES’ denotes the predictions filtered by our QES metric.
CIML on SPG CIML on SDG
Method P R F IoU Method P R F IoU
Nonzero .159 .987 .273 .079 DMVN* [26] .553 .357 .434 .276
OSTU [29] .812 .589 .683 .497 DMVN [26] .438 .568 .495 .317
DMVN* [26] .329 .618 .430 .276 DMAC* [51] .583 .538 .559 .410
DMVN [26] .682 .781 .728 .578 DMAC [51] .703 .622 .660 .518
DMAC* [51] .617 .624 .620 .432 SACM [14] (Conf.) .778 .819 .798 .702
DMAC [51] .746 .710 .728 .573 Corr-DINO (Ours) .816 .862 .838 .744
DASS (Ours) .860 .921 .889 .835 Corr-DINO (QES) .939 .971 .954 .912
TABLE IX: Ablation study of CIML on SPG. ’Forged’ denotes using the forged image as input, ’Auth.’ denotes using the authentic image as input, ’Diff.’ denotes using the absolute difference map between the forged and authentic image as input.
Num. Backbone Forged Auth. Diff. P R F1 IoU
(1) VGG [52] $\times$ $\times$ .423 .365 .392 .236
(2) VGG [52] $\times$ .800 .857 .827 .718
(3) VGG [52] $\times$ $\times$ .771 .857 .812 .741
(4) VGG [52] .825 .899 .860 .781
(5) VAN [30] $\times$ $\times$ .573 .533 .552 .408
(6) VAN [30] .860 .921 .889 .835
VII-C Experiments on the CAAAv2 Models (CIML task)
This subsection evaluates the effectiveness of the proposed CAAAv2 in the CIML task by comparing its DASS and Corr-DINO models against other CIML models on SPG and SDG test datasets, respectively. We also conduct detailed ablation studies for in-depth analysis.
Implementation Details. Given that real-world forgeries in the IMD20 [32] dataset closely resemble the target images for annotation, we use a portion of this dataset to evaluate model performance using IoU and binary F1-score metrics. We categorize the forged images in IMD20 into SPG or SDG and randomly split them into training and test sets with a 3:1 ratio. CASIAv2 [31] and approximately 1M images synthesized from the COCO [49] dataset are also used for training. Input images are resized to $\times512$, and a consistent training pipeline is applied across all methods.
Comparison Study. Table VIII presents comparison results for SPG (left) and SDG (right) separately. Our method clearly outperforms these previous methods, primarily because our category-aware paradigm significantly reduces task difficulty. As shown in the right side of Table VIII, our Corr-DINO, prior to QES filtering, achieves an IoU of 0.744. This is 4.2 points higher than the SACM [14] (our conference version), demonstrating that our frozen-denoising paradigm notably alleviates the overfitting issue when training on limited data. Notably, DMVN [26] and DMAC [51] models trained on both SPG and SDG data perform worse than those trained exclusively on one category. This is attributed to the confusion caused by similar backgrounds in SPG image pairs.
Ablation Study for the DASS Model Input. Table IX presents the results, confirming that both the image pair and the image difference map are essential for accurate prediction in SPG. The image difference map between forged and authentic images can roughly indicate the forged regions, while the images themselves provide semantic information to help the model denoise the difference map.
Ablation Study for the Corr-DINO Model Backbone. The results are shown in Table X, settings (1-19). For the ConvNeXt CNN model, the trainable version performs better. For all ViT models, however, frozen versions perform significantly better. Since ViT models tend to overfit with limited training data, freezing the backbone alleviates catastrophic forgetting and overfitting. LoRA [53] fine-tuned versions exhibit similar performance to their frozen counterparts, as LoRA also mitigates forgetting. The frozen DINOv2-Base [13] backbone performs optimally and is therefore adopted.
Ablation Study of the Corr-DINO Model Modules. The results are shown in the bottom of Table X, setting (20) is the baseline without any proposed modules, setting (24) is the full model with all proposed modules. In settings (21, 22, 23), we remove the proposed Multi-Aspect Denoiser, Feature Super Resolution and Learnable Aggregation modules from the full model (24) respectively. Setting (24) significantly outperforms (21). This validates that correlation features extracted from the frozen DINOv2 are highly noisy, and that denoising through multi-aspect feature integration is essential for accurate predictions. Setting (24) also significantly outperforms (22). This confirms that initial correlation features are too coarse for precise mask predictions, and our Feature Super Resolution module effectively addresses this by reconstructing fine-grained details. Finally, Setting (24) is superior to (23) because our learnable aggregation mitigates information loss inherent in rule-based aggregation.
Ablation Study of the QES. The last two rows of Table X show the results. The results of setting (25) are obtained by evaluating only the predictions from setting (24) where QES>>0.5. QES-filtered predictions from setting (25) achieve an IoU of 0.912, representing a 16.8 point increase over setting (24). This confirms that QES effectively filters out unsatisfactory predictions, thereby notably improving annotation quality. Table XI confirms QES’s high relevance to the IoU metric without relying on ground truth for evaluation.
Our Models are Adequate Auto-Annotators. In Table VIII, both our DASS and Corr-DINO achieve high IoU scores. Considering the errors in IMD20’s ground truth [54], especially in SPG images, our actual accuracy is higher. Clearly, our methods are sufficient for obtaining accurate auto-annotations.
TABLE X: Ablation study of CIML on SDG. ‘Freeze‘ denotes freezing the backbone and ‘L‘ denotes LoRA [53] fine-tuning. ‘MAD‘ denotes the Multi-Aspect Denoiser. ‘FSR‘ denotes the Feature Super Resolution. ‘LA‘ denotes the Learnable Aggregation. ‘QES‘ denotes filtering the prediction with our QES.
Num. Backbone Freeze MAD FSR LA QES P R F1 IoU
(1) ConvNeXt-B [50] $\times$ $\times$ .740 .690 .714 .595
(2) ConvNeXt-B [50] $\times$ .635 .554 .592 .414
(3) Eff.SAM [55] $\times$ $\times$ .483 .624 .544 .381
(4) Eff.SAM [55] $\times$ .606 .502 .549 .403
(5) CLIP-L [56] $\times$ $\times$ .665 .509 .577 .453
(6) CLIP-L [56] L $\times$ .744 .736 .740 .587
(7) CLIP-L [56] $\times$ .749 .721 .735 .596
(8) DINOv1-B [57] $\times$ $\times$ .693 .512 .589 .456
(9) DINOv1-B [57] L $\times$ .693 .742 .717 .595
(10) DINOv1-B [57] $\times$ .700 .700 .700 .586
(11) DINOv2-S [13] $\times$ $\times$ .739 .653 .693 .584
(12) DINOv2-S [13] L $\times$ .783 .733 .758 .637
(13) DINOv2-S [13] $\times$ .777 .781 .772 .667
(14) DINOv2-B [13] $\times$ $\times$ .705 .574 .633 .472
(15) DINOv2-B [13] L $\times$ .794 .861 .826 .727
(16) DINOv2-B [13] $\times$ .816 .862 .838 .744
(17) DINOv2-L [13] $\times$ $\times$ .667 .517 .582 .454
(18) DINOv2-L [13] L $\times$ .803 .843 .823 .725
(19) DINOv2-L [13] $\times$ .809 .850 .829 .732
(20) DINOv2-B [13] $\times$ $\times$ $\times$ $\times$ .644 .605 .624 .469
(21) DINOv2-B [13] $\times$ $\times$ .699 .666 .682 .526
(22) DINOv2-B [13] $\times$ $\times$ .822 .728 .772 .664
(23) DINOv2-B [13] $\times$ $\times$ .804 .838 .821 .719
(24) DINOv2-B [13] $\times$ .816 .862 .838 .744
(25) DINOv2-B [13] .939 .971 .954 .912
TABLE XI: Ablation study on the threshold of QES.
QES Threshold 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
kept sample ratio 2.45 1.90 1.84 1.72 1.51 1.00 0.73 0.50 0.31 0.06
IMD20 SDG IoU .744 .825 .842 .865 .889 .912 .924 .957 .967 .968
TABLE XII: Image manipulation localization performance on SACP [58] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [7] and CASIAv2 [31] datasets for model pretraining
Num. Methods Pretraining data IoU F1
(1) DFCN [59] Image-Net [60] .466 .607
(2) MVSS-Net [20] Image-Net [60] .401 .534
(3) PSCC-Net [37] Image-Net [60] .482 .620
(4) RRU-Net [43] Image-Net [60] .517 .651
(5) CFL-Net [61] Image-Net [60] .433 .571
(6) DTD [62] Image-Net [60] .588 .712
(7) TIFDM [63] Image-Net [60] .576 .703
(8) Web-IML (Ours) Image-Net [60] .773 .862
(9) Web-IML (Ours) ADE20k [64] .774 .863
(10) Web-IML (Ours) Forgery Baseline .778 .865
(11) Web-IML (Ours) (10) + DEFACTO [8] .767 .856
(12) Web-IML (Ours) (10) + PSCC Synthetic [37] .762 .853
(13) Web-IML (Ours) (10) + AIGC Synthetic [48] .756 .851
(14) Web-IML (Ours) (10) + MIMLv1 [14] (Conference) .792 .872
(15) Web-IML (Ours) (10) + Web Supervision (Ours) .800 .880
VII-D Experiments on the Downstream Document IML Task
This subsection evaluates the generalization of the Web-IML model and the universal applicability of our web-supervision method on an important downstream task: document IML.
We fine-tune our Web-IML under various pretraining settings on the handcrafted document IML benchmark SACP [58], using the same training and test configurations as prior works [63]. Our Web-IML notably outperforms prior methods. Moreover, Table XII shows that settings (11, 12, 13), which incorporate additional training data from DEFACTO [8], PSCC [37], and OpenSDID [48] into the baseline (Setting 10), perform worse than the baseline. This confirms that simply increasing the scale and diversity of pretraining data does not lead to better generalization. In contrast, the model pretrained on our web supervision data (15) shows a significant 2.2 point IoU improvement over the baseline (10), and is 0.8 points higher than the model pretrained on MIMLv1 (14) from our conference paper. These results demonstrate that the cheap data from our web supervision method is of sufficient quality to enable models to learn semantic-agnostic forensic features and generalize across unseen real-world IML tasks.
VIII Conclusion
We propose CAAAv2, a novel category-aware framework that solves the constrained image manipulation localization problem by separately handling Shared Donor Group (SDG) and Shared Probe Group (SPG) image pairs through Difference-Aware Semantic Segmentation and Correlation DINO with frozen DINOv2 backbone, dramatically reducing overfitting while enabling high-quality automatic annotation. With CAAAv2 and a novel metric QES, we auto-annotate manually forged images from the web and harvest a large-scale, diverse, high-quality dataset MIMLv2, comprising 246,212 manually forged images with pixel-level annotations. This dataset considerably addresses the data scarcity issue in IML. Furthermore, we propose Object Jitter, an effective data generation method that leverages authentic web images to improve model performance in challenging cases. To better leverage web-scale supervision, we also introduce Web-IML, a new model for IML. Experiments confirm the strong generalization ability of our method, which achieves unprecedented performance gains and surpasses previous state-of-the-art methods by significant margins. We believe this work establishes a scalable framework for continuously expanding training data as new manually forged images emerge online. Its universal applicability is demonstrated across multiple benchmarks. This research transforms image forensics, shifting reliance from limited, expensive handcrafted datasets to abundant, continuously growing web resources, thus paving the way for more robust and generalizable IML systems.
Figure 12: Qualitative comparison on image manipulation localization.

### v2

@@
 […]
-Index Terms:
-Image manipulation localization, automatic annotation, web supervision, image forensics.
+Figure 1: Left: Definition and characteristics of the Shared Donor Group (SDG) and Shared Probe Group (SPG).
+Right: In contrast to previous works, our CAAAv2 adopts category-aware and prior-feature-denoising paradigm that notably reduces task difficulty and overfitting.
+We also propose using the CAAAv2 models to auto-annotate web-scale manually forged images.
+This approach significantly addresses the data scarcity issue for IML and notably improves the generalization of all IML models.
+footnotetext: This research is supported in part by NSFC (Grant No.: 62476093) and GD-NSF (No.2021A1515011870).
+(Corresponding author: Lianwen Jin.) Chenfan Qu, Huiguo He, and Lianwen Jin are with the School of Electronics and Information Engineering, South China University of Technology, Guangdong 510641, China.
+(Email: 202221012612@mail.scut.edu.cn, hehuiguo@scut.edu.cn, eelwjin@scut.edu.cn) Yiwu Zhong is with the School of Intelligence Science and Technology, Peking University, Beijing 100091, China.
+(Email: yiwu-zhong@outlook.com) Bin Li is with the College of Electronics and Information Engineering, Shenzhen University, Shenzhen 518060, China.
+(Email: libin@szu.edu.cn)
 I Introduction
-The misuse of manipulated images can lead to fraud and the spread of misinformation, posing serious threats to social media security [1].
-To this end, Image Manipulation Localization (IML) has gained increasing attention in recent years as a critical research problem [2, 3].
+The misuse of manipulated images can lead to fraud and the spread of misinformation, posing serious threats to social media security [1, 2, 3, 4, 5, 6, 7, 8].
+To this end, Image Manipulation Localization (IML) has garnered increasing attention in recent years [9, 10, 11, 12, 13, 14, 15, 16].
 A central challenge in IML lies in the scarcity of large-scale high-quality data [17], which is essential for preventing model overfitting.
-Generating such data is extremely labor-intensive and time-consuming [5, 6].
-It requires not only crafting convincing manipulated images but also providing precise pixel-level annotations of the forged regions.
-Synthetic data has been explored as an alternative [7, 8], but the resulting images often differ substantially from real-world forgeries, creating a significant domain gap that hinders model generalization [4].
-Recent methods have attempted to close this gap by leveraging noise domain features [9, 10] and semantic suppression techniques [6], showing progress in improving robustness.
-Nevertheless, these methods do not fully resolve the fundamental issue: the lack of large-scale, manually forged data.
-As a result, overfitting remains a persistent challenge in IML.
-Figure 1: When an image is tampered with by copy-pasting foreground objects from original images, the resulting original-tampered image pair belongs to Shared Donor Group (SDG).
-When an image is forged by directly modifying elements in place, the pair belongs to Shared Probe Group (SPG).
-Figure 2: We propose a novel CIML method that processes the SPG and SDG images separately.
-It is used for automatic annotation, constructing a large-scale, high-quality dataset.
-The dataset notably enhances the generalization of image manipulation localization models.
-Additionally, the Object Jitter method is introduced to further enlarge the dataset size and data diversity.
-Meanwhile, there are abundant manually forged images publicly available on the web, often accompanied by their corresponding originals.
-We propose to leverage such resources to address the data scarcity problem in IML.
-Specifically, we propose leveraging Constrained Image Manipulation Localization (CIML) models to automatically generate pixel-level mask annotations for these unlabeled forged images.
-Our key insight is that CIML simplifies the manipulation localization task by comparing a forged image with its corresponding original.
-This reduction in task complexity allows CIML to produce more accurate and reliable annotations for real-world manipulations.
-As a result, CIML offers a practical solution for automatically generating large-scale, high-quality training datasets for IML.
-However, despite recent progress on less challenging data, existing CIML methods remain inadequate as reliable automatic annotators for complex real-world images, due to three major limitations.
-First, most prior approaches rely on a single correlation-based model to handle all input data [11, 12], which we argue is a suboptimal design.
-In practice, image pairs consisting of forged images and their corresponding original images can be categorized into two distinct types based on their relationship (see Fig.
-1): (1) Shared Donor Group (SDG), where forged image is created by copying objects from the original image and pasting them onto another image.
-In SDG, the shared content between the real and fake images is the foreground object.
-(2) Shared Probe Group (SPG), where forged image is created by directly modifying the original image.
-In SPG, the shared content is the less object-specific background.
-Although correlation-based methods work reasonably well on SDG data, they are less effective on SPG data.
-This is because the shared backgrounds in SPG cover larger areas and offer fewer discriminative features compared to the localized, distinctive foregrounds in SDG.
-Training a single correlation model on both groups can therefore lead to confusion, resulting in degraded generalization and performance.
-Second, existing CIML methods overlook an important signal, that is the difference maps obtained by subtracting forged images from their originals.
-These difference maps consistently highlight manipulated regions and offer strong localization cues on SPG data.
-However, previous studies did not treat SPG and SDG images separately, thereby leaving this valuable signal underutilized.
-Third, CIML on SDG is inherently much more challenging than on SPG.
-Unlike SPG, CIML on SDG cannot benefit from image difference information.
-Additionally, SDG exhibits greater visual complexity and diverse manipulation patterns.
-Unfortunately, high-quality training data for SDG is extremely limited, making it difficult to train robust and generalizable models for real-world scenarios.
-To tackle these problems, we propose a novel paradigm termed as Category-Aware Auto-Annotation v2 (CAAAv2), which treats image pairs in SDG and SPG separately.
-Our CAAAv2 comprises three components.
-At the beginning, a classifier is employed to determine whether an input image pair belongs to SDG or SPG.
-This classifier can be trained effectively through self-supervised learning using unlabeled images.
-Moving forward, a Difference-Aware Semantic Segmentation model that utilizes both the image pairs and their difference maps for accurate constrained manipulation localization in SPG.
-Additionally, a Correlation DINO model that improves generalization on SDG with a frozen DINOv2 [13] encoder, a Learnable Aggregation module, a Feature Super Resolution module and a Multi-Aspect Denoiser.
-The proposed CAAAv2 is an upgraded version of our CAAA at CVPR 2024 [14].
-In our earlier work, overfitting occurred in SDG because a trainable backbone was used during training with limited data.
-To address this issue, we now employ a frozen ViT backbone, which has resulted in significant improvements for CIML on SDG.
+Creating such data is extremely labor-intensive and time-consuming [18], requiring not only the crafting of convincing forgeries but also the precise pixel-level annotation of manipulated regions.
+Although alternatives like synthetic data [19, 20] or AI-generated forgeries [21] have been explored, they often exhibit significant discrepancies from real-world manipulations, thereby limiting model generalization to practical scenarios [17].
+Meanwhile, numerous unlabeled, manually forged images are publicly available over the Internet, often accompanied by their corresponding originals.
+To address the scarcity of annotated data, we propose to leverage these unlabeled forgeries by automatically generating pixel-level mask annotations using Constrained Image Manipulation Localization (CIML) models.
+Our key insight is that CIML can reduce the complexity of manipulation localization by comparing the forged image with its original, resulting in more accurate and reliable annotations.
+Therefore, CIML offers a practical solution for automatically generating large-scale, high-quality training datasets for IML.
+However, despite recent progress with CIML on simple synthetic images, existing methods are unreliable for complex, real-world forgeries due to two major limitations:
+First, previous methods employ a suboptimal paradigm that relies on a single correlation-based model for all input types [22, 23].
+In practice, image pairs consisting of forged images and their original counterparts can be categorized into two distinct groups (see Fig.
+1): (1) Shared Donor Group (SDG) denotes image pairs that share foreground, with forged images generated by copying objects from the original image; (2) Shared Probe Group (SPG) denotes image pairs that share background, where forged images are obtained by modifying the original image in-place.
+While correlation-based methods perform reasonably well on SDG data, they falter on SPG data because the shared region (the background) often lacks distinctive features.
+Forcing a single model to handle both groups leads to confusion and degraded performance.
+Second, the CIML task itself is challenging and also suffers from a scarcity of high-quality training data.
+Previous works provide the model with only the image pair, relying entirely on the network to learn the complex correspondence between the images in each pair.
+Without clear and explicit guidance, models struggle to learn this correspondence effectively and are prone to overfitting on limited training data.
+To overcome these challenges, we propose a novel paradigm: Category-Aware Auto-Annotation v2 (CAAAv2). It achieves superior generalization via two key innovations:
+(1) Category awareness.
+We treat image pairs in SDG and SPG separately through a classifier at start.
+This separation avoids the confusion of joint training and allows us to deploy distinct, optimized pipelines that leverage the unique correspondence of each category.
+This classifier can be effectively trained on unlabeled images.
+(2) Task transformation.
+We reframe the challenging CIML problem as a far simpler denoising task.
+This is achieved by introducing novel prior features that provide strong initial guidance for localization.
+Crucially, these prior features are extracted via zero-shot frozen functions that are resistant to overfitting.
+For SDG pairs, where the forged regions are typically shared objects, we extract features from both images using a frozen DINOv2 encoder and compute their correlation map.
+The resulting map, named DINO correlation, serves as a prior that roughly highlights the manipulated regions (Fig.
+1).
+The model’s task is reduced to simply denoising this prior to produce a precise mask.
+For SPG pairs, where the forged regions are the difference between the images, we obtain the prior map by subtracting the forged image from its original.
+Through task transformation, we significantly reduce task difficulty and bridge the gap between synthetic and real-world data, leading to substantially better generalization.
 Subsequently, we collect a large amount of manually forged images from the Internet and then annotate their forged regions with the proposed CAAAv2.
 This approach considerably alleviates the data scarcity in image manipulation localization.
-To ensure the reliability of the annotations, we propose Quality Evaluation Score (QES), a novel metric that automatically evaluates annotation quality and filters out inadequate ones.
-Experiments demonstrate that our MIMLv2 dataset significantly improves generalization of various models on widely used benchmarks, as shown in Fig.
-2.
-Furthermore, the enhanced precision of CAAAv2 ensures that a greater number of high-quality samples are retained after filtering.
-Consequently, MIMLv2 is both larger and more diverse than our conference version MIMLv1 [14], better supporting the development of IML models.
-To further enhance the scale and diversity of web supervision, we propose Object Jitter.
-This method generates semantically reasonable high-quality artifacts by introducing slight distortions to random objects, such as enlargement, overexposure, or texture alteration.
-Then, we train IML models to localize these distorted objects.
-Our approach effectively supplements MIMLv2 with additional authentic web images.
-This also differs from our conference paper that only uses MIMLv1 dataset to improve IML models.
-To best leverage our high-quality web-scale supervision for IML, we further propose a new model named Web-IML.
-It comprises an encoder, a Multi-Scale Perception module, and a Self-Rectification module.
-The Multi-Scale Perception module enables the model to better integrate information from multiple perspectives.
-The Self-Rectification module allows the model to learn to identify and correct its mistakes.
+To ensure the reliability of the annotations, we propose Quality Evaluation Score (QES), a novel metric that automatically assesses and filters out inadequate annotations.
+While this pipeline yields a high-quality dataset, we observed a relative lack of sophisticated copy-move forgeries that feature only subtle edge artifacts.
+Direct copy-move synthesis is not effective since it brings obvious semantic artifacts.
+To address this, we introduce Object Jitter, a data synthesis method that simulates these challenging cases by applying minor perturbations (e.g., enlargement) to random objects on authentic images, effectively supplementing our dataset.
+In our CVPR 2024 paper, we propose to address the data scarcity of IML by auto-annotating images from the web with a category-aware approach CAAA.
+However, CAAA did not incorporate a prior-feature-denoising paradigm for its SDG branch, leading to overfitting on limited training data.
+In this paper, we propose CAAAv2, an improved version of our CAAA that resolves this issue by leveraging frozen DINO-based prior features.
+The enhanced accuracy of CAAAv2 ensures that a greater number of samples are retained after filtering.
+Consequently, the collected dataset MIMLv2 is both larger and more diverse than our conference version MIMLv1 [24], better supporting the development of IML models (Figure 1).
+Moreover, different from our conference paper that only uses MIMLv1 dataset to improve IML models, our Object Jitter approach further effectively supplements MIMLv2 with high-quality hard samples.
+To better leverage our high-quality web-scale supervision for IML, we further propose a new model named Web-IML, which comprises a Multi-Scale Perception module and a Self-Rectification module.
+The former enables the model to better integrate information from multiple perspectives; the latter allows the model to learn to identify and correct its mistakes.
 Our Web-IML significantly outperforms previous methods on various widely used benchmarks, being smaller yet more effective than the APSC-Net in our conference version.
 In summary, our main contributions are as follows:
-  • We propose to address the challenging IML task through the abundant web forgeries and their high-quality annotations distilled from a less challenging task, CIML.
-  • We introduce a novel method CAAAv2 for CIML, along with a novel metric QES designed to assess prediction quality.
-Together, they are used to automatically annotate manually forged web images, constructing a large-scale, diverse, high-quality dataset MIMLv2.
+  • We propose to address the scarcity of IML training data by leveraging the less challenging CIML task to auto-annotate forged images collected from the web.
+  • We introduce CAAAv2, a novel CIML paradigm that significantly enhances generalization.
+This is achieved through a category-aware design, which treats distinct image groups separately, and a prior-denoising approach, which reframes CIML as a simpler denoising task.
+  • With CAAAv2 and a novel metric QES, we automatically annotate manually forged web images, constructing a large-scale, diverse, high-quality dataset MIMLv2.
 This dataset considerably alleviates the data scarcity in IML.
-  • To further increase the size and diversity of the training data, we introduce Object Jitter, which generates high-quality artifacts using authentic web images.
-  • We design a new model Web-IML, to better leverage the web-scale supervision. Extensive experiments have confirmed the effectiveness of the proposed method.
+  • We introduce Object Jitter, which generates high-quality, challenging-to-localize artifacts, effectively supplementing our dataset.
+To better leverage the web-scale supervision, we design a novel model Web-IML.
+Extensive experiments have validated its effectiveness.
 Figure 2: The overall pipeline of our proposed CAAAv2 paradigm.
-The input original-manipulated image pairs are first fed into a classifier to be classified into SPG or SDG.
-Then, we use Difference Aware Semantic Segmentation and Correlation DINO to predict forged regions in SPG and SDG respectively.
+The input original-forged image pairs are first categorized as either SDG or SPG with a classifier.
+Then, we use Corr-DINO and Difference-Aware Semantic Segmentation to process SDG and SPG pairs respectively.
+Both models first extract a prior feature that roughly highlights forgeries and then denoise it.
 II Related Works
-Image Manipulation Localization aims to identify forged regions within images [15, 16, 17, 18, 19] .
-Due to costly data collection, existing handcrafted datasets are tiny in size, leading to overfitting in many models [6, 4].
-To address overfitting, some studies incorporated handcrafted features.
-RGB-N [9] utilized SRM [9] filters to suppress semantic bias.
+II-A Image Manipulation Localization (IML)
+Image Manipulation Localization aims to identify forged regions within images [25, 26, 27, 28, 29, 30, 31, 32, 33].
+The scale of existing handcrafted datasets is limited due to the high cost of manual collection and annotation, inevitably leading to overfitting [18, 17].
+To mitigate this issue, some studies have incorporated handcrafted features.
+For instance, RGB-N [34] applied SRM [34] filters to suppress semantic bias.
 MVSS-Net [35] and MGQFormer [36] further exploited learnable noise domain modeling.
 CAT-Net [19], ObjectFormer [37] and HiFi-Net [38] leveraged frequency features to localize manipulations.
-However, the handcrafted features are mostly noisy and unstable, hindering further improvements in model performance.
-Recently, pure vision models with semantic suppression techniques have achieved better performance [24, 25].
+However, the handcrafted features are noisy and unstable, hindering further improvements in model performance.
+Recently, pure vision models have made notable progress [39, 1].
 NCL [17] employed contrastive learning to improve generalization.
 SparseViT [18] utilized dilated attention to reduce semantic bias.
-Despite the progress made, these works still suffer a lot from over-fitting due to scarce high-quality data.
-To this end, we propose to address data scarcity by constructing a large-scale high-quality dataset, with the Internet images accurately annotated in an automatic way.
-Constrained Image Manipulation Localization aims at identifying the forged image region with reference to its original image [26].
-Most prior works relied on correlation matching and treated SDG and SPG image pairs uniformly.
-DMVN [26] computed correlation maps to localize similar objects across image pairs.
-DMAC [11] incorporated dilated convolution for enhanced spatial information.
-AttentionDM [12] employed attention mechanism for better performance.
-MSTAF [27] performed correlation in both the encoder and decoder to extract better features.
-While these methods achieved notable progress on less challenging datasets (e.g., synthetic COCO [26]), their performance is limited in real-world scenarios characterized by high resolution, large variations, and complex details.
-In contrast to previous approaches, our method treats SPG and SDG separately, leveraging the differences between forged images and their originals.
-We are also the first to employ a frozen backbone for this task, effectively reducing overfitting.
+PIM-Net [1] explicitly modeled the super-pixel inconsistency to alleviate biased learning.
+Despite these advances, overfitting remains a significant challenge due to the fundamental issue: the scarcity of high-quality training data.
+To address this fundamental issue, we propose to construct a large-scale, high-quality dataset by automatically annotating Internet images with a less challenging task, CIML.
+II-B Constrained Image Manipulation Localization (CIML)
+Constrained Image Manipulation Localization aims to identify forged regions in an image by leveraging its original counterpart [40].
+Previous works relied on correlation matching and handled SDG and SPG image pairs uniformly.
+For instance, DMVN [40] computed correlation maps to localize similar objects across image pairs.
+DMAC [22] incorporated dilated convolution to enhance spatial information.
+AttentionDM [23] employed an attention mechanism to improve performance.
+MSTAF [41] performed correlation in both the encoder and decoder to extract richer features.
+While these methods achieved notable progress on less challenging datasets (e.g., synthetic COCO [40]), their performance is limited in real-world forgeries, which often involve high-resolution images, significant variations, and complex details.
+In contrast, our approach reduces task difficulty and addresses overfitting by (1) explicitly harnessing the distinct characteristics of SPG and SDG pairs; and (2) transforming the CIML problem into a denoising task through prior features extracted from frozen functions.
 III Category-Aware Auto-Annotation v2
-For constrained image manipulation localization (CIML), previous works didn’t consider the critical discrepancy between SPG and SDG images, and processed them using a single correlation-based model.
-We argue that such a paradigm is sub-optimal and the reasons are as follows:
-First, the similar regions in SDG images are typically foreground objects (e.g., the cats in the SDG branch of Fig.
-3) that have specific, similar shapes and distinctive features.
-In contrast, the similar regions in SPG images are background, which generally lacks distinctive features for accurate correlation matching (e.g., for the images of the SPG branch in Fig.
-3, a patch of snow in the background is similar to all other snow patches).
-These background regions will cause confusion for correlation-based models, especially in complex scenarios.
-Second, the difference between the forged image and its original in an SPG pair serves as an important cue.
-Most areas in an SPG image pair are almost identical and spatially aligned (e.g., the image pair in the SPG branch of Fig.
-3).
-Simply subtracting between them yields a difference map that can always highlight the manipulated regions.
-However, this information is difficult to utilize in correlation-matching-based models and was therefore overlooked in prior works.
-Based on these observations, we propose a novel paradigm for the CIML task, Category-Aware Auto-Annotation v2 (CAAAv2).
-The key idea is to process SPG and SDG images independently, as shown in Fig.
-3.
-First, a classifier (described in Section III.A) categorizes the input image pairs as either SPG or SDG.
-For SPG pairs, we apply Difference-Aware Semantic Segmentation (Section III.B).
-For SDG pairs, we process the images using Correlation DINO (Section III.C).
-Importantly, the trained models are further used to auto-annotate unlabeled forged images, thereby addressing the data scarcity issue in IML tasks.
-The difference between CAAAv2 and its conference version CAAA lies in the SDG branch model.
-In CAAA, we processed SDG image pairs using a trainable CNN structure, which tended to overfit the limited training data.
-In contrast, CAAAv2 employs a well pre-trained frozen ViT backbone, thereby reducing overfitting to task-specific data.
-Moreover, the ViT backbone naturally integrates global information, thereby alleviating semantic misalignment caused by scale diversity.
-III-A Self-Supervised Classification
-To achieve the classification of SPG and SDG, we propose training a classifier via self-supervised learning on unlabeled images.
-Given an image, we apply random augmentations (e.g.
-color jitter, image compression) and manipulations (copy-paste, splicing or removal).
+III-A Motivation and Framework Overview
+For CIML, previous works overlooked the critical discrepancy and characteristics between SPG and SDG images, treating them using a single fully-tunable correlation model.
+We argue that this paradigm is suboptimal for two reasons:
+1.
+Conflicting Correlation Signals.
+The nature of the ”shared region” differs drastically between the two categories.
+In SDG pairs, the shared region is salient foreground objects with distinctive features (e.g., the lion in SDG of Fig.
+2), providing a strong signal for correlation matching.
+Conversely, in SPG pairs, the shared region is the background, which lacks the unique features required for reliable matching (e.g., the sky in SPG of Fig.
+2).
+Forcing a single correlation model to simultaneously handle both groups creates a conflicting objective that degrades performance.
+2.
+Lack of Explicit Guidance.
+Learning to comprehend and fully leverage the correspondence between original and manipulated images is non-trivial.
+With scarce training data, CIML models tend to find shortcuts to minimize training loss, often overfitting to either forgery-agnostic semantics in small-scale handcrafted datasets or monotonous patterns in synthetic data.
+Without explicit guidance on how to utilize the correspondence within an image pair, models often fail to generalize to diverse, real-world manipulations.
+To overcome these challenges, we propose Category-Aware Auto-Annotation v2 (CAAAv2), a novel paradigm for CIML.
+Our core idea is to process SDG and SPG images independently, while explicitly integrating image correspondence as a guiding prior feature.
+Specifically, as shown in Fig.
+2, CAAAv2 introduces a classifier to categorize the input as either SPG or SDG, and then applies a category-specific prior to guide localization.
+For SDG pairs, we introduce DINO Correlation, a prior that captures shared foreground objects by correlating dense features from a frozen DINOv2 encoder.
+For SPG pairs, a simple Difference Map effectively identifies manipulated regions by subtracting the two images.
+These priors embed domain-specific knowledge about forgery patterns into strong, zero-shot signals.
+By generating these priors with frozen, non-trainable functions, we reframe the complex task of open-ended forgery localization into a simpler, more constrained denoising problem.
+Here, the model’s objective is to refine the coarse prior into a precise mask, using the input images as context.
+This paradigm shift significantly reduces task difficulty and minimizes the generalization gap between synthetic and real-world forgeries, leading to substantial performance improvements.
+Importantly, the trained models are further used to auto-annotate large-scale, unlabeled manual forgeries from the web, thereby addressing the data scarcity issue in IML.
+Whereas CAAA overfits on SDG data, CAAAv2 provides explicit guidance by introducing DINO Correlation as a strong prior.
+This prior is generated using a frozen DINOv2, whose global receptive field naturally aligns multi-scale semantics, yielding substantial gains in real-world generalization.
+III-B Self-Supervised Classification
+Training a classifier to distinguish between SDG and SPG is a non-trivial task due to the lack of labeled data.
+To address this challenge, we propose training a classifier via self-supervised learning on unlabeled images.
+Given an image, we apply random augmentations (e.g., color jitter, image compression) and in-place manipulations (copy-paste, splicing, or removal).
 The manipulated image and the original image form an SPG pair.
@@
 For example, in Fig.
-8, the two images in any SPG pair are nearly identical and spatially aligned, whereas images in any SDG pair are distinctly different.
+6, the two images in any SPG pair are nearly identical and spatially aligned, whereas images in any SDG pair are distinctly different.
 These characteristics are universally applicable and strong enough.
 Therefore, this classification task is quite simple, enabling accurate separation of images into the two groups.
-Figure 4: Manipulated images often degrade during transmission.
-As a result, the absolute difference between a tampered image and its original does not accurately indicate the forged region.
-Our method effectively denoises the difference map by leveraging semantic information.
-III-B Difference-Aware Semantic Segmentation (DASS)
-Ideally, for an SPG image pair, the absolute difference between the authentic and forged images would directly reveal the forged region.
-However, due to transmission noise and image degradation [28], the absolute difference map is often imprecise.
-As shown in Fig.
-4, almost every area in the difference map shows nonzero values because of noise.
-Even after binarizing the difference map using an OTSU [29] adaptive threshold, noisy highlights persist, especially in high-frequency regions such as edges.
-To address this issue, we propose denoising the difference map by incorporating semantic information from the images.
-Specifically, we concatenate the authentic image, the forged image, and their absolute difference map along the channel dimension, and feed the result into a semantic segmentation model.
-Our segmentation model employs a VAN [30] encoder for semantic feature extraction and utilizes the same Multi-Aspect Denoiser as in our Corr-DINO to denoise the difference map.
-III-C Correlation DINO (Corr-DINO)
-CIML on SDG requires detecting shared objects at various scales within complex backgrounds.
-This task is considerably more challenging than SPG, where targets are already highlighted through image difference maps.
-As a result, synthetic or small-scale handcrafted datasets are largely insufficient for training a robust model for complex real-world scenarios.
-To tackle the overfitting issue, we introduce a novel frozen-denoising paradigm (see Fig.
-5).
-Our key insight is that overfitting can be mitigated by employing a frozen encoder and subsequently denoising its output features.
-Building on this paradigm, we developed the Corr-DINO model (see Fig.
-6), which consists of five components:
+III-C Corr-DINO
+For Shared Donor Group (SDG) pairs, the manipulated regions are mostly the objects copied from the original image.
+The localization task therefore reduces to identifying these shared foreground objects.
+To exploit this correspondence, we introduce a novel prior feature, termed DINO Correlation.
+This prior is generated by extracting dense features from both images using a frozen DINOv2 encoder and then computing their cross-correlation.
+The resulting map provides a strong yet coarse signal that highlights the shared objects, which often are the forged regions.
+The task is thus simplified into refining this coarse prior to a precise segmentation mask.
+Based on this, we propose the Corr-DINO model.
+As illustrated in Figure 3, it consists of four components:
 1.
-Frozen Encoder for feature extraction.
-We extract features from both the tampered image $I_{a}$ and its original image $I_{b}$ with a frozen DINOv2 [13] backbone.
-This ViT backbone has learned generalized representations through self-supervised learning on vast-scale web images.
-The features from the last ViT layer, denoted as $F_{a1}$ and $F_{b1}$, contain rich semantic information and are used to compute correlation features.
-2.
-Correlation module for correlation features calculation.
+Prior Feature Extractor for prior feature extraction.
+We extract object features from both the tampered image $I_{a}$ and its original image $I_{b}$ with a frozen DINOv2 [42] backbone, which has learned generalized representations through self-supervised learning on vast-scale web images.
+The features from the last DINO layer, denoted as $F_{a1}$ and $F_{b1}$, contain rich semantic information and are used to compute correlation features $F_{a\_corr}$ and $F_{b\_corr}$.
 $F_{a\_corr}=[Corr(F_{a1},F_{b1}),Corr(F_{a1},F_{a1})]$, $F_{b\_corr}=[Corr(F_{b1},F_{a1}),Corr(F_{b1},F_{b1})]$.
 The same correlation function $Corr$ as most previous works [23, 41, 24] is adopted.
-The detailed $Corr$ function is presented in the Appendix.
-Figure 5: Different from previous works that employ learnable encoders and rule-based feature aggregation for constrained image manipulation localization on SDG image pairs, we propose a novel frozen-denoising paradigm.
-This paradigm offers significantly better generalization in real-world scenarios.
-Figure 6: The overall pipeline of our proposed Correlation DINO.
+Figure 3: The overall pipeline of our proposed Corr-DINO.
 First, we employ a frozen DINO backbone to extract image features.
@@
 Green lines indicate operations used only during training with synthetic data, due to the lack of mask labels for the real images.
-3.
-A Learnable Aggregation module for dimension reduction.
+2.
+Learnable Aggregation module for dimension reduction.
 Previous works simply select the top-K activated channels with the highest average values in $F_{a\_corr}$ and $F_{b\_corr}$ to reduce dimension.
@@
 To overcome this, we propose a learnable aggregation module that uses trainable parameters to reduce the channel count to K.
-Following prior methods, we also include statistics such as average and maximum values across the channel dimension.
+We also include statistics of average and maximum values across the channel dimension.
 $F_{a\_aggr}=[C_{1\times 1}(R(C_{1\times 1}(F_{a\_corr}))),Avg(F_{a\_corr}),Max(F_{a\_corr})]$.
-Here, $C_{1\times 1}$ is $1\times 1$ conv-layer and $R$ is ReLU.
+Here, $C_{1\times 1}$ is $1\times 1$ conv-layer, $R$ is ReLU.
+$Avg$ and $Max$ are average and maximum pooling respectively.
 A similar process is used for $F_{b\_corr}$ to generate $F_{b\_aggr}$.
-4.
+3.
 Feature Super Resolution module for image detail reconstruction.
-The features from the last ViT layer are rich in semantics but lack fine image details [13], making $F_{a\_aggr}$ and $F_{b\_aggr}$ too coarse for fine-grained mask predictions.
+The features from the last DINOv2 layer are rich in semantics but lack fine image details [42], making $F_{a\_aggr}$ and $F_{b\_aggr}$ too coarse for fine-grained mask predictions.
 To reconstruct image details, we leverage features from the last 2-4 layers of DINO, which are known to be more sensitive to edges and fine details [42].
@@
 We then fuse them with $1\times 1$ conv-layers.
-$F_{asr\_i}=Conv_{1\times 1}(Conv_{1\times 1}(F_{a})+Conv_{1\times 1}(F_{aggr\_ai}))$ for iin(1,2,3,4)i\,in\,(1,2,3,4).
 A similar process is applied to $F_{b\_aggr}$.
-5.
-Multi-Aspect Denoiser module for features denoising.
-Since the ViT encoder is frozen, its output features inevitably contain significant task-irrelevant noise.
+4.
+Multi-Aspect Denoiser for feature denoising.
+Since the DINOv2 encoder is frozen, its output features inevitably contain significant task-irrelevant noise.
 To achieve precise mask prediction, we propose denoising the correlation features by integrating multi-perspective information.
 This module fuses multi-scale features $F_{asr\_1,2,3,4}$ in a top-down manner.
-$F^{\prime}4=Conv_{3\times 3}([F_{asr\_4},Avg(F_{asr\_4})])$, $F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{asr\_n}))$ for iin(1,2,3)i\,in\,(1,2,3).
-Here, $Conv_{3\times 3}$ and $Conv_{1\times 1}$ are conv-layers with kernel size 3 and 1, and $Avg$ is global average pooling.
-The fused features are further concatenated and denoised using a set of dilated $3\times 3$} conv-layers.
-The output is concatenated to $F^{\prime}1$ and processed by a $3\times 3$} conv-layer to yield the mask prediction.
-Our Correlation DINO is optimized end-to-end using cross-entropy loss. It significantly alleviates the overfitting problem in challenging real-world CIML tasks for SDG.
+The fused features are further concatenated and denoised using a set of dilated conv-layers.
+The output is concatenated to $F^{\prime}1$, which is used to yield the mask prediction.
+More details are provided in the Appendix.
+Figure 4: Manipulated images often degrade during transmission.
+As a result, the absolute difference between a tampered image and its original does not accurately indicate the forged region.
+Our method effectively denoises the difference map by leveraging semantic information.
+III-D Difference-Aware Semantic Segmentation (DASS)
+SPG pairs are characterized by in-place manipulation, where the manipulated image is generated by directly altering its original.
+This process inherently ensures precise spatial alignment and a shared background, meaning that the forged regions are always highlighted by the image difference.
+Leveraging this strong correspondence, we use the difference map as a prior feature to guide model learning.
+Given a manipulated image $I_{m}$ and its original $I_{o}$, the absolute difference map $D_{mo}=abs(I_{m}-I_{o})$ can roughly highlight forged regions.
+However, transmission noise and image degradation often contaminate the difference map, making it too noisy for precise localization.
+As illustrated in Fig.
+4, nearly every area exhibits nonzero values due to noise, and even adaptive binarization [43] fails to denoise.
+To address this issue, we propose denoising the difference map by incorporating semantic information from the images.
+Specifically, we concatenate $I_{m}$, $I_{o}$, and $D_{mo}$ along the channel dimension, and feed them into a semantic segmentation model.
+This model integrates a VAN [44] encoder for extracting rich semantic features and employs the same Multi-Aspect Denoiser (III-C) as in our Corr-DINO for denoising the difference map.
 IV MIMLv2 Dataset
 In this section, we introduce MIMLv2, a large-scale, diverse, and high-quality dataset.
-The key idea is to leverage constrained image manipulation localization models trained on existing datasets to automatically annotate manually forged images collected from the web.
-To ensure high annotation quality, we propose a novel metric that filters out inadequate annotations.
+The key idea is to leverage CIML models trained on existing datasets to auto-annotate manual forgeries collected from the web, thereby addressing the data scarcity in IML.
+To ensure high annotation quality, we propose a novel metric that filters out inadequate annotations without requiring ground-truth for evaluation.
 Our MIMLv2 is annotated using our CAAAv2, which is significantly more accurate than its conference version, CAAA [24].
@@
 Compared to MIMLv1 [24], MIMLv2 offers greater scale and diversity, making it a more effective resource for training IML models.
-Figure 7: The overall construction pipeline of our MIMLv2 dataset.
+Figure 5: The construction pipeline of our MIMLv2 dataset.
 IV-A Dataset Construction
 As illustrated in Fig. 5, we construct MIMLv2 through the following steps:
+1.
 Image Collection.
@@
 On this website, the images are manually manipulated and thus have high-quality, diverse forged regions.
-Additionally, we exclude images that already appear in the evaluation datasets (Section VII) and remove duplicates by checking their MD5 and pHash.
+Additionally, we exclude damaged, duplicated images and those that already appear in the evaluation datasets [45, 46, 47, 48, 49, 50, 51] with MD5 [52] and pHash [53].
+2.
 Auto-Annotation.
 We utilize the CAAAv2 models, as proposed in Section III, to obtain mask annotations for the collected images.
-These models are sufficiently trained with one million synthetic COCO samples [26], along with the handcrafted datasets CASIAv2 [31] and IMD20 [32].
+The models are effectively trained on synthetic COCO [40], CASIAv2 [46] and IMD20 [45] datasets.
 […]
-Figure 8: Some images and their corresponding mask annotations from the proposed MIMLv2 dataset. Left: SPG image pairs. Right: SDG image pairs.
+Figure 6: Some images and their corresponding mask annotations from the proposed MIMLv2 dataset.
 IV-B Dataset Highlights
-We present a few samples of the proposed MIMLv2 dataset in Fig. 8. The main highlights of our dataset are as follows:
+We present a few samples of our MIMLv2 dataset in Fig.
+6.
+The forged regions are precisely annotated with high-quality masks, demonstrating the effectiveness of our auto-annotation approach.
+The main highlights of our dataset include:
   • High quality.
-Image manipulations in the proposed dataset are meticulously crafted by humans.
-Such data can teach models to detect real-world forgery, rather than merely overfitting to simple patterns found in synthetic data.
+The proposed dataset contains high-quality image manipulations crafted by humans.
+Such data is crucial to training models that can detect real-world forgeries rather than overfit to synthetic data.
   • Large Scale.
-As shown in Table I, the proposed dataset comprises 246,212 manually forged images.
-This is dozens of times more samples than previous handcrafted IML datasets (e.g., $\approx$120 times more samples than IMD20 [32]).
-  • Board Diversity.
-MIMLv2 comprises images of various sizes, styles, and manipulation types (e.g., copy-move, splicing, removal, AIGC).
-These images are created by hundreds of thousands of individuals using various software.
-Such broad diversity can considerably enhance the generalization ability of deep IML models.
-  • Modern Style.
-MIMLv2 contains abundant modern images that were recently captured and forged, reflecting current photography and forgery technologies.
+The proposed dataset is substantially larger than existing handcrafted IML datasets.
+As shown in Table I, it comprises 246,212 manually forged images, which is about 120 times more than IMD20 [45].
+  • Broad Diversity.
+MIMLv2 comprises images of various sizes, styles, and manipulation types (e.g., copy-move, splicing, removal, AIGC), created by more than 10k individuals and various software.
+This diversity significantly enhances the generalization of IML models.
+  • Updated Forgeries.
+MIMLv2 contains numerous modern images that were recently captured and forged, reflecting current photography and forgery technologies.
 In contrast, the CASIA dataset [46] proposed over a decade ago, features blurred small images and outdated forgeries.
-Therefore, MIMLv2 better meets the current requirements for IML.
+Therefore, our MIMLv2 better suits current IML requirements.
   • Strong Scalability.
@@
 CASIAv1 [46] 2013 800 921 (246, 384)-(500, 334)
-CASIAv2 [31] 2013 7461 5,123 (160, 240)-(901, 600)
+CASIAv2 [46] 2013 7,461 5,123 (160, 240)-(901, 600)
 Coverage [47] 2016 100 100 (190, 334)-(472, 752)
@@
 In Wild [55] 2018 0 201 (650, 650)-(2736, 3648)
+MISD [51] 2021 618 300 (160, 240)-(901, 600)
 IMD20 [45] 2020 414 2,010 (193, 260)-(4437, 2958)
+CIMD [50] 2024 102 102 (2048, 1365)-(2048, 1365)
 MIMLv1 [24] 2024 0 123,150 (45, 120)-(13846, 9200)
@@
 V Object Jitter
-We experimentally find that the IML model performance can be further improved with more high-quality data.
-To this end, we introduce Object Jitter, an effective method that enhances IML models by generating high-quality image artifacts online during training from authentic web images.
-Previous works synthesize tampered images by randomly cropping and pasting image objects [37, 9].
-However, these data synthesis methods often produce overly obvious artifacts and semantically unreasonable content, as shown on the left in Fig.
-9.
-Models trained on such data typically overfit to simple cases and perform poorly on real-world forgeries, where artifacts are subtle and tampered content is semantically reasonable.
-DEFACTO [8] attempts to ensure semantic reasonableness by designing elaborate rules.
-It manually identifies eight object types with consistently similar shapes and appearances (e.g., the road sign in the middle of Fig.
-9); copy-move or splicing operations are restricted to objects of the same type.
-Because DEFACTO applies only to a limited number of object classes, the generated data lacks diversity, causing models to overfit the semantics of these specific object types.
-Furthermore, DEFACTO heavily relies on accurate object class annotations, which significantly limits its broader application.
-To overcome these challenges, we propose a novel idea: instead of altering image semantics through tampering, we simulate subtle artifacts.
-Building on this idea, we introduce Object Jitter, a novel method detailed in Fig 10.
-Given an input image, we readily obtain a segmentation mask for a random object using off-the-shelf segmentation models like SAM [38].
-The selected object is then processed using one or a combination of the following three random operations:
-  • Size Jitter. Slightly enlarge the selected object’s region and its corresponding mask by a random scale.
-  • Exposure Jitter. Slightly overexpose the selected object.
-  • Texture Jitter. Slightly alter the texture of the selected object region by applying a random combination of JPEG compression, reverse JPEG compression [39], and blur.
-After processing, we add a random smooth transition on the edges to minimize visual artifacts.
-We apply Object Jitter to 1 to 3 random objects within an input image.
-The mask regions of these processed objects are recorded and subsequently considered as tampered regions during training.
-Our method successfully addresses existing issues: First, it always preserves the semantics of the input image, thereby preventing models from overfitting to simple cases with abnormal semantics.
-Second, by applying subtle changes to objects that already naturally blend into the background, the generated visual image artifacts are unobvious.
-Third, our method is unconstrained and universally applicable.
-Consequently, Object Jitter consistently produces high-quality artifacts and effectively supplements our MIMLv2 dataset.
-Figure 9: Comparison between different data generation methods. Our method is universally applicable and always produces semantically reasonable forgeries.
+Sophisticated copy-move forgeries, where an object is copied and pasted within the same image, are difficult to detect.
+They introduce only subtle edge artifacts while preserving source texture and noise patterns.
+Their scarcity in web-sourced data leads models to rely more on common noise and compression anomalies, causing them to fail on these hard cases.
+Existing synthesis methods, designed to generate training data for such forgeries, are flawed: random copy-paste techniques [56, 34] often create semantically incoherent artifacts (e.g., a train on food, Fig.
+7), while more constrained approaches like DEFACTO limit tampering to a few objects of consistent shape, lacking object diversity.
+Both issues can cause models to overfit to unrealistic cues.
+To address this gap, we propose Object Jitter, a novel strategy that simulates these forgeries by subtly distorting existing objects.
+The core operation is Size Jitter: we slightly enlarge a randomly selected object, covering its original footprint.
+This mimics a realistic copy-move forgery by preserving semantics and texture while introducing only faint boundary artifacts.
+As illustrated in Fig.
+9, the process involves (1) segmenting an object using ground truth or SAM [57], (2) scaling the object and its mask by a small random factor, and (3) applying a smooth alpha-blended transition at the new boundary to minimize extraneous visual cues.
+To further enhance data diversity, we introduce two additional minor perturbations: Exposure Jitter (slight overexposure) and Texture Jitter (subtle modifications via JPEG compression, reverse JPEG [58], and blur operations).
+By preserving semantic plausibility and being applicable to any object, Object Jitter provides a source of challenging training examples.
+This compels the model to learn the subtle boundary artifacts characteristic of sophisticated forgeries, thereby improving its robustness and generalization.
+Figure 7: Comparison between different data generation methods.
 Figure 8: The proposed Object Jitter method.
 Given a randomly selected object from an authentic image, we slightly enlarge or overexpose it, or alter its texture.
-The processed object is regarded as tampered object for training.
+The processed object is regarded as a tampered object for training.
 VI Web-IML
@@
 As illustrated in Fig.
-11, Web-IML comprises a feature extractor, a Multi-Scale Perception module to integrate information from multiple perspectives, and a Self-Rectification module that enables the model to identify and correct its mistakes.
-Figure 11: The overall framework of the proposed Web-IML model.
-Multi-Scale Perception module motivation.
-During meticulous image forensic analysis, humans often zoom in and out the image repeatedly, comparing between diverse observations to assist their final prediction.
-To mimic the human perception way, we introduce the Multi-Scale Perception module.
-Multi-Scale Perception module pipeline.
-Specifically, given four feature maps extracted from the backbone model $F_{1},F_{2},F_{3},F_{4}$}, we first extract global features in different scales: $G_{i}=Conv_{1\times 1}(Avg_{i}(F_{4}))\,for\,i\,in\,(1,2,3,6)$, where $Conv_{1\times 1}$ denotes conv-layer with kernel size 1 and $Avg_{i}$ denotes average pooling function with output size $i\times i$.
-The global features are then resized, concatenated to $F_{4}$ and fused, dimension reduced to get $F^{\prime}4$.
-$F^{\prime}4=Conv_{3\times 3}([G_{1},G_{2},G_{3},G_{6},F_{4}])$.
+9, Web-IML comprises a feature extractor, a Multi-Scale Perception module to integrate information from multiple perspectives, and a Self-Rectification module that enables the model to identify and correct its mistakes.
+Figure 9: The overall framework of the proposed Web-IML.
+VI-A Multi-Scale Perception Module
+Motivation.
+This module is designed to mimic the meticulous image forensic analysis, where humans often zoom in and out the image repeatedly, comparing diverse observations to assist their final prediction.
+Pipeline.
+Given the feature maps $F_{1},F_{2},F_{3},F_{4}$} extracted from the encoder, we first extract global features of different scales through multi-scale average pooling.
+The global features are then fused and dimension-reduced to get $F^{\prime}4$.
 Next, the high-level features are fused with the low-level features recursively.
-$F^{\prime}_{n}=Conv_{3\times 3}(F^{\prime}_{n+1}+Conv_{1\times 1}(F_{n}))\,for\,i\,in\,(1,2,3)$.
 This process fully integrates high-level semantics with low-level image details, resulting in the multi-scale features $F^{\prime}_{1},F^{\prime}_{2},F^{\prime}_{3},F^{\prime}_{4}$}.
-Self-Rectification module motivation.
-Due to the challenging nature of the image manipulation localization task, forensic models inevitably overlook key information and make mistakes.
-To minimize negligence, we propose the Self-Rectification module.
-With this module, our model can learn to check and correct its initial predictions.
-Self-Rectification module pipeline.
+More details are provided in the Appendix.
+VI-B Self-Rectification Module
+Motivation.
+This module is designed to make the model learn to check and correct its initial predictions, thereby minimizing the overlooking of key information and mistakes.
+The Self-Rectification (SR) module improves upon the Self-Calibration (SC) module in our conference paper from two aspects: First, in SC, we map the initial prediction mask to a convolution kernel, use it to convolve the mask and multiply the result by the multi-scale features.
+Despite its effectiveness, learning the map function to a convolution kernel is difficult, and directly multiplying the convolved mask often causes information loss in the black mask region.
+In our SR, we map the initial prediction mask to a feature map, and concatenate it with the multi-scale features, addressing the issues.
+Second, we introduce a Nested Channel Attention module that enables more meticulous analyses of the rectification feature.
+Pipeline.
 As shown at the bottom of Fig.
-11, we obtain the initial mask prediction $P_{n}$ from $F^{\prime}_{1}$ via a $\times1$ conv-layer.
+9, we obtain the initial mask prediction $P_{n}$ from $F^{\prime}_{1}$.
 We then adaptively obtain a rectification feature $F_{r}$ through a tiny Rectification Feature Mapping module ($RFM$), which consists of 5 cascaded conv-layers.
-$F_{r}=RFM(P_{n})$.
 The $F_{r}$ informs the model of suspected tampered regions and potentially incorrect prediction regions.
-Then $F_{r}$ is fused with $F_{a}$ with a $\times3$ conv-layer to get $F_{c}$.
-$F_{c}=Conv_{3\times 3}([F_{a},P_{c}])$.
-To enable in-depth, meticulously analyses of $F_{c}$, we introduce the Nested Channel Attention module.
-As shown in the top left of Fig.
-11, it can be formulated as $G_{1}=Avg(F_{c})W_{1}$, $G_{2}=G_{1}W_{2}$, $A_{2}=\sigma(W_{3}G_{2})$, $A_{1}=\sigma(W_{4}G_{1}*A_{2})$, $F_{o}=F_{c}*A_{1}$.
-Here, $G_{1,2}$ are global vectors, $A_{1,2}$ are attention maps, $W_{1,2,3,4}$ are learnable parameters, $Avg$ is global average pooling and $\sigma$ is the sigmoid function.
-The output $F_{o}$ is further processed by four $\times3$ conv-layers with dilation sizes $1,2,3,6$ respectively, the output features are concatenated and fused with a $\times3$ conv-layers.
-The mask prediction $P_{f}$ is obtained from $F_{o}$ with a $\times1$ conv-layer.
-To get the refined final prediction $P_{f}$, we replace $P_{n}$ with $P_{f}$, replace $F_{a}$ with $F_{o}$, then repeat the self rectification once again.
-Our Web-IML is optimized with the cross entropy loss calculated between the prediction mask and the ground-truth.
+$F_{r}$ is fused with $F_{a}$ to get $F_{c}$.
+Then, in the Nested Channel Attention module, we obtain the global vectors $G_{1}$ from $F_{c}$, and subsequently higher-level global vectors $G_{2}$ from $G_{1}$.
+Then $G_{1}$ is weighted by the attention map calculated by $G_{2}$ and the results are further used to calculate the channel attention weights for $F_{c}$.
+The output $F_{o}$ is further processed by four dilated conv-layers, the output features are concatenated and fused to produce the mask prediction $P_{f}$.
+To get the refined final prediction $P_{f}$, we replace $P_{n}$ with $P_{f}$, replace $F_{a}$ with $F_{o}$, then repeat the self-rectification once more.
+Our Web-IML is optimized with cross entropy loss in an end-to-end manner.
 TABLE II: Comparison on image manipulation localization.
-”WS” is our proposed web supervision, which trains models using data from the MIMLv2 and Object Jitter.
-”Avg*” is the average score excluding IMD20.
-”Param.” is the number of model parameters.
+‘WS‘ is our proposed web supervision, which trains models using data from the MIMLv2 and Object Jitter.
+‘Avg*‘ is the average score excluding IMD20.
+‘Param.‘ is the number of model parameters.
+CAT-Net, TruFor and SparseViT were trained with the IMD20 dataset and therefore not tested on it.
 Method CASIAv1 [46] Coverage [47] NIST16 [48] IMD20 [45] CocoGlide [49] CIMD [50] MISD [51] Avg.* Average Param.
@@
 TruFor [49] .630 .692 .446 .522 .279 .348 - - .294 .362 .322 .418 .426 .549 .400 .482 - - 67M
+SparseViT [18] .768 .819 .456 .504 .322 .381 - - .327 .389 .257 .322 .421 .538 .425 .492 - - 50M
 PIM [1] .512 .566 .188 .251 .225 .280 .340 .419 .327 .404 .177 .229 .389 .519 .303 .375 .308 .381 168M
@@
 SparseViT [18] .888 .884 .869 .881 .877 .886 .881
+NCL [17] .912 .856 .831 .840 .806 .843 .819
 ERMPC [64] .895 .893 .877 .892 .871 .894 .888
@@
 Web-IML (Ours) .942 .935 .904 .914 .900 .940 .909
-TABLE V: Ablation study on our Web Supervision method with Web-IML model.
+TABLE V: Ablation study on our Web Supervision method.
 ‘PSCC-Synthetic‘ denotes the synthetic data from the PSCC-Net [56].
@@
 Num. Training Data IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
-(1) Baseline .823 .857 .508 .551 .412 .465 .370 .422 .285 .339 .629 .673 .430 .545 .494 (+0.0%) .550 (+0.0%)
+(1) Web-IML Baseline .823 .857 .508 .551 .412 .465 .370 .422 .285 .339 .629 .673 .430 .545 .494 .550
 (2) +PSCC-Synthetic [56] .813 .845 .503 .546 .348 .392 .361 .410 .177 .211 .598 .654 .330 .441 .447 (-9.5%) .500 (-9.1%)
@@
 (4) +AIGC-Synthetic [21] .817 .849 .273 .296 .220 .276 .284 .337 .394 .462 .287 .340 .373 .429 .376 (-24.%) .420 (-23.%)
+(5) +WSCL [66] .815 .849 .467 .510 .415 .468 .411 .464 .293 .347 .630 .672 .431 .544 .495 (+0.2%) .551 (+0.2%)
 (6) +MIMLv1 [24] (Conf.) .821 .854 .561 .616 .520 .587 .692 .768 .275 .325 .789 .845 .496 .610 .594 (+20.%) .658 (+20.%)
@@
 TABLE VI: Ablation study on our Object Jitter method.
-‘COCO‘ denotes using the COCO [49] dataset images for Object Jitter.
-’SA’ denotes using 400k random images from the SA-1B [38] dataset for Object Jitter.
+‘COCO‘ denotes using the COCO [67] images for Object Jitter.
+‘SA‘ denotes using 400k random images from the SA-1B [57] dataset for Object Jitter.
 ‘C.G.‘ denotes the CocoGlide dataset [49].
-Num Object Jitter Operation
-Image
-Source
-CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] C.G. [10] CIMD [40] MIFD [41] Average
+Num Object Jitter Operation Image Source CASIAv1 [46] Coverage [47] NIST16 [48] IMD20 [45] C.G. [49] CIMD [50] MISD [51] Average
 Size Texture Exposure IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
@@
 (7) SA .846 .879 .715 .768 .591 .670 .698 .776 .439 .507 .759 .821 .496 .613 .649 .719
+TABLE VII: Comparison study of our web-supervision on IML task.
+Blue number for improvement and Red number for decrease.
+‘B.L.‘ denotes baseline setting.
+‘PSCC-Synth.‘ denotes the synthetic dataset of PSCC-Net [56], ‘AIGC-Synth.‘ denotes the AIGC forgery dataset OpenSDID [21].
+‘O.J.‘ denotes our Object Jitter method.
+Training Data CASIAv1 [46] Coverage [47] NIST16 [48] IMD20 [45] CIMD [50] Average
+IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$
+PSCC-Net B.L. .401 0 .430 0 .197 0 .218 0 .247 0 .295 0 .125 0 .156 0 .161 0 .221 0 .226 0 .264 0
++PSCC-Synth. .352 12% .379 12% .147 25% .159 27% .230 7% .274 7% .101 19% .125 20% .101 37% .124 44% .186 18% .212 20%
++DEFACTO[20] .394 2% .429 1% .231 17% .256 17% .223 10% .270 8% .157 26% .171 10% .120 25% .159 28% .225 0% .257 3%
++AIGC-Synth. .346 16% .377 14% .109 81% .123 77% .091 171% .142 108% .110 14% .144 8% .124 30% .170 30% .156 45% .191 38%
++WSCL [66] .439 9% .464 8% .201 2% .220 1% .253 2% .300 2% .192 54% .235 51% .177 10% .239 8% .252 12% .293 11%
++MIMLv1 (Conf.) .609 52% .649 51% .395 101% .477 119% .402 63% .476 61% .470 276% .541 247% .206 28% .251 14% .416 84% .479 81%
++MIMLv2 .602 50% .642 49% .402 104% .498 128% .405 64% .485 64% .575 360% .662 324% .354 120% .483 119% .468 107% .554 110%
++MIMLv2+O.J .609 52% .651 51% .451 129% .550 152% .406 64% .486 65% .570 356% .658 322% .335 108% .464 110% .474 110% .562 113%
+CAT-Net B.L. .660 0 .703 0 .245 0 .286 0 .239 0 .287 0 .157 0 .192 0 .229 0 .263 0 .306 0 .346 0
++PSCC-Synth. .410 38% .437 38% .215 12% .232 19% .288 21% .331 15% .121 23% .148 23% .160 30% .189 28% .239 22% .268 23%
++DEFACTO[20] .673 2% .715 2% .200 18% .230 20% .220 8% .261 9% .164 4% .200 4% .130 43% .156 41% .277 9% .312 10%
++AIGC-Synth. .646 2% .685 3% .136 80% .147 95% .170 41% .242 19% .130 21% .165 16% .088 160% .117 125% .234 31% .271 28%
++WSCL [66] .668 1% .710 1% .240 2% .281 2% .245 3% .292 2% .204 30% .246 29% .238 4% .273 4% .319 4% .361 4%
++MIMLv1 (Conf.) .691 5% .728 4% .302 23% .389 36% .353 48% .422 47% .547 248% .629 228% .397 73% .477 81% .458 50% .529 53%
++MIMLv2 .684 4% .720 2% .368 50% .443 55% .338 41% .416 45% .562 258% .644 235% .470 105% .548 108% .484 58% .554 60%
++MIMLv2+O.J .703 7% .742 6% .355 45% .437 53% .358 50% .440 53% .575 266% .656 242% .497 117% .575 119% .498 63% .570 65%
+TruFor B.L. .755 0 .794 0 .363 0 .411 0 .320 0 .377 0 .248 0 .289 0 .296 0 .346 0 .396 0 .443 0
++PSCC-Synth. .750 1% .798 1% .340 7% .384 7% .325 2% .373 1% .253 2% .298 3% .278 6% .331 5% .389 2% .437 2%
++DEFACTO[20] .752 0% .791 0% .322 13% .357 15% .214 50% .260 45% .209 19% .247 17% .184 61% .217 59% .336 18% .374 18%
++AIGC-Synth. .746 1% .791 0% .220 65% .261 57% .213 50% .263 43% .221 12% .267 8% .121 146% .156 122% .304 30% .348 28%
++WSCL [66] .758 1% .797 1% .362 0% .410 0% .326 1% .373 1% .307 24% .352 22% .310 5% .359 4% .412 4% .458 3%
++MIMLv1 (Conf.) .755 0% .794 0% .385 6% .458 10% .352 9% .408 8% .596 158% .678 157% .340 13% .401 14% .486 18% .548 19%
++MIMLv2 .767 2% .794 0% .387 6% .458 10% .407 21% .472 20% .618 160% .696 158% .370 20% .422 18% .510 22% .568 22%
++MIMLv2+O.J .770 2% .811 2% .466 22% .511 20% .439 27% .502 25% .619 160% .700 159% .309 4% .357 3% .521 24% .576 23%
+SparseViT B.L. .702 0 .747 0 .320 0 .352 0 .288 0 .343 0 .199 0 .239 0 .177 0 .232 0 .337 0 .383 0
++PSCC-Synth. .627 11% .677 9% .304 5% .333 5% .256 11% .311 9% .196 1% .243 1% .169 5% .232 0% .310 8% .359 6%
++DEFACTO[20] .692 1% .730 2% .257 20% .283 20% .310 8% .364 6% .186 7% .225 6% .155 12% .198 15% .320 5% .362 6%
++AIGC-Synth. .533 24% .583 22% .119 63% .141 60% .133 54% .174 49% .088 56% .110 54% .033 81% .045 81% .181 46% .211 45%
++WSCL [66] .709 1% .755 1% .324 1% .356 1% .301 4% .352 3% .267 34% .314 31% .184 4% .241 4% .357 6% .404 6%
++MIMLv1 (Conf.) .714 2% .762 2% .394 23% .458 30% .306 6% .359 5% .535 169% .617 158% .215 21% .270 16% .433 28% .493 29%
++MIMLv2 .723 3% .771 3% .441 38% .509 45% .338 17% .396 15% .544 173% .624 161% .234 32% .289 25% .456 35% .518 35%
++MIMLv2+O.J .720 3% .767 3% .477 49% .547 55% .346 20% .409 19% .551 178% .633 165% .252 42% .325 40% .469 40% .536 40%
+IMDPrompter B.L. .718 0 .765 0 .463 0 .522 0 .346 0 .419 0 .261 0 .310 0 .354 0 .414 0 .428 0 .486 0
++PSCC-Synth. .682 5% .729 5% .473 2% .531 2% .295 15% .356 15% .262 0% .312 1% .286 19% .337 19% .400 7% .453 7%
++DEFACTO[20] .701 2% .749 2% .439 5% .508 3% .284 18% .347 17% .243 7% .286 8% .281 21% .329 21% .390 9% .444 9%
++AIGC-Synth. .630 12% .678 11% .254 45% .289 45% .218 37% .263 37% .222 15% .264 15% .263 26% .309 25% .317 26% .361 26%
++WSCL [66] .701 2% .750 2% .485 5% .543 4% .360 4% .433 3% .285 9% .336 8% .356 1% .417 1% .437 2% .496 2%
++MIMLv1 (Conf.) .724 1% .769 1% .542 17% .610 17% .418 21% .501 20% .601 130% .696 125% .392 11% .456 10% .535 25% .606 25%
++MIMLv2 .729 2% .773 1% .564 22% .635 22% .426 23% .512 22% .614 135% .708 128% .418 18% .490 18% .550 29% .624 28%
++MIMLv2+O.J .735 2% .778 2% .626 35% .706 35% .433 25% .521 24% .635 143% .721 133% .425 20% .498 20% .571 33% .645 33%
 VII Experiments
@@
 The model is trained for 200k iterations with a batch size of 16, using a $\times512$ input size, consistent with previous works [18, 49].
-We employ Cross-Entropy loss and AdamW [14] optimizer, with the learning rate linearly decaying from 1e-4 to 1e-6.
+We employ cross entropy loss and AdamW [69] optimizer, with the learning rate linearly decaying from 1e-4 to 1e-6.
 CASIAv2 [46] and tampCOCO [19] serve as baseline training sets.
@@
 Comparison Study.
-Table II shows that our Web-IML significantly outperforms existing methods on multiple real-world image forgery benchmarks.
-For example, it achieves an average IoU that is 24.1 points higher than TruFor.
+Table II shows that our Web-IML significantly outperforms existing methods.
+For example, it achieves an average IoU that is 21.6 points higher than SparseViT.
 Furthermore, Web-IML consistently outperforms our conference version, APSC-Net, across various settings.
@@
 Qualitative comparison is illustrated in Fig.
-12.
+10.
 Ablation Study.
@@
 This indicates that the proposed MP module effectively integrates high-level and low-level features to detect diverse image artifacts.
-Setting (3), which further incorporates the Self-Rectification (SR) module, yields an additional 3.9 point IoU improvement over Setting (2), as SR enables the model to detect and correct its errors.
+Setting (3), which further incorporates the Self-Rectification (SR) module, yields an additional 3.9 points IoU improvement over Setting (2), as SR enables the model to detect and correct its errors.
 Setting (4) outperforms (3) via our nested attention module that enables in-depth analysis.
@@
 Our web-supervision encompasses both the MIMLv2 dataset and the Object Jitter method.
-We compare the performance of the same IML model trained on data created by our web-supervision approach against other synthetic methods.
+We compare the performance of the same IML model trained on data created by our web-supervision approach against other synthetic methods and WSCL [66], a representative weakly supervised framework.
 Implementation Details.
@@
 When supplementing the baseline training sets with additional IML training sets, the total training volume is kept constant with a uniform sampling ratio across all datasets.
+For models integrated with WSCL, training involved two types of data: direct training with the fully (pixel-level) annotated tampCOCO [19] and CASIAv2 [46] datasets, and WSCL-specific training with weakly (image-level) annotated MIMLv2 and SA [57] dataset images.
+The training volume and other configurations remain the same as ours.
 Comparison Study for the MIMLv2 Dataset.
 Table V presents the IML performance of the Web-IML model when trained on different datasets.
-The baseline (1) uses the two commonest training sets: tampCOCO [7] and CASIAv2 [31].
+The baseline (1) uses the two most common training sets: tampCOCO [19] and CASIAv2 [46].
 Incorporating synthetic training data from PSCC-Net [56], DEFACTO [20] or OpenSDID [21] diminishes the average model performance (settings 2, 3, 4).
-This is because the first two synthetic datasets do not notably surpass the baseline data in quality or diversity, while the AIGC forgeries from OpenSDID differ significantly from non-AIGC forgeries.
-These results confirm that merely increasing training data scale does not lead to improved performance.
-In contrast, including the MIMLv1 (setting 5) or MIMLv2 (setting 6) datasets significantly improves model performance by an average of 10 and 12 points, respectively, without increasing training or inference burden.
+This is because the first two synthetic datasets do not notably surpass the baseline data in quality or diversity, while the AIGC forgeries from OpenSDID [21] differ significantly from non-AIGC forgeries.
+These results confirm that simply increasing the scale of training data does not lead to improved performance.
+In contrast, including the MIMLv1 (setting 6) or MIMLv2 (setting 7) datasets significantly improves model performance by an average of 10 and 12 points, respectively, without increasing training or inference burden.
+Additionally, although WSCL (setting 5) offers improvements over baselines, its effectiveness is considerably lower than that of our web-supervision.
+This is because our method explicitly produces accurate mask annotations, by leveraging the image correspondence that substantially reduces task difficulty.
+Conversely, WSCL lacks the ability to leverage this valuable information, and relies heavily on unstable noise domain features.
 These results confirm that our MIMLv1 and MIMLv2 datasets considerably alleviate data scarcity and overfitting issues in real-world scenarios.
 This is attributable to their large scale, manual forgery, and considerably higher quality and diversity compared to other datasets.
-Notably, MIMLv2 yields greater improvement due to its larger quantity of high-quality samples and superior diversity compared to MIMLv1.
+Notably, MIMLv2 yields greater improvement due to its larger quantity and superior diversity compared to MIMLv1.
 Comparison Study for the Object Jitter.
-Table V shows that including the training data generated by our Object Jitter method further improves the model when combined with MIMLv1 (setting 7 outperforms 5) or MIMLv2 (setting 9 outperforms 6) datasets, effectively supplementing them.
-This is attributed to our Object Jitter method’s consistent generation of diverse, high-quality samples, which further enriches the training data.
+Table V shows that including the training data generated by our Object Jitter method further improves the model when combined with MIMLv1 (setting 8 outperforms 6) or MIMLv2 (setting 10 outperforms 7) datasets.
+This improvement is attributed to Object Jitter’s consistent generation of high-quality hard samples, which effectively supplements the training data by compensating for the relative scarcity of copy-move forgeries.
 Ablation Study for the Object Jitter.
@@
 Since most Internet forgeries are not generated by copy-move, they exhibit exposure and texture artifacts.
-The size jitter operation, without altering the object’s exposure or texture, reduces model’s reliance on exposure or texture cues, thus better supplementing MIMLv2.
-Furthermore, Setting (6) applies the full Object Jitter method using COCO [49] as the image source, instead of 400k random images from SA [38] in setting (7).
+The size jitter operation, without altering the object’s exposure or texture, reduces the model’s reliance on exposure or texture cues, thus better supplementing MIMLv2.
+Furthermore, Setting (6) applies the full Object Jitter method using COCO [67] as the image source, and setting (7) uses 400k random images from SA [57].
 The similar performance of Settings (6) and (7) demonstrates that the Object Jitter method is robust to the choice of image source.
-Conclusions Verification on Other IML Models.
+Verification of Conclusions on Other IML Models.
 We extended our verification process to other IML models.
-Specifically, we evaluated MIMLv2 and Object Jitter on PSCC-Net [37], CAT-Net [7], and TruFor [10] using their official model code.
-Table VII presents the results, which clearly indicate that MIMLv1, MIMLv2, and Object Jitter all lead to significantly higher scores.
+Specifically, we evaluated MIMLv2 and Object Jitter on PSCC-Net [56], CAT-Net [19], TruFor [49], SparseViT [18] and IMDPrompter [70] (a large model based on SAM [57]) using the same training data and configuration as ours.
+The results in Table VII clearly indicate that MIMLv1, MIMLv2, and Object Jitter all lead to significantly higher scores.
 MIMLv2, in particular, consistently yields significantly greater improvement than MIMLv1 (conference version).
@@
 These results robustly confirm our previous analyses and conclusions.
-TABLE VII: Comparison study of our web-supervision on IML task.
-Blue number for improvement and Red number for decrease.
-’B.L.’ denotes baseline setting using tampCOCO [7] and CASIAv2 [31] datasets as training data.
-’PSCC-Synth.’ denotes the synthetic dataset of PSCC-Net [37], ’AIGC-Synth.’ denotes the AIGC forgery dataset OpenSDID [48].
-’O.J.’ denotes our Object Jitter method.
-Training Data CASIAv1 [31] Coverage [34] NIST16 [35] IMD20 [32] CIMD [40] Average
-IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$ IoU $\Delta$ F1 $\Delta$
-PSCC-Net B.L. .401 0 .430 0 .197 0 .218 0 .247 0 .295 0 .125 0 .156 0 .161 0 .221 0 .226 0 .264 0
-+PSCC-Synth. .352 12% .379 12% .147 25% .159 27% .230 7% .274 7% .101 19% .125 20% .101 37% .124 44% .186 18% .212 20%
-+DEFACTO[8] .394 2% .429 1% .231 17% .256 17% .223 10% .270 8% .157 26% .171 10% .120 25% .159 28% .225 0% .257 3%
-+AIGC-Synth. .346 16% .377 14% .109 81% .123 77% .091 171% .142 108% .110 14% .144 8% .124 30% .170 30% .156 45% .191 38%
-+MIMLv1 (Conf.) .609 52% .649 51% .395 101% .477 119% .402 63% .476 61% .470 276% .541 247% .206 28% .251 14% .416 84% .479 81%
-+MIMLv2 .602 50% .642 49% .402 104% .498 128% .405 64% .485 64% .575 360% .662 324% .354 120% .483 119% .468 +107% .554 110%
-+MIMLv2+O.J .609 52% .651 51% .451 129% .550 152% .406 64% .486 65% .570 356% .658 322% .335 108% .464 110% .474 110% .562 113%
-CAT-Net B.L. .660 0 .703 0 .245 0 .286 0 .239 0 .287 0 .157 0 .192 0 .229 0 .263 0 .306 0 .346 0
-+PSCC-Synth. .410 38% .437 38% .215 12% .232 19% .288 21% .331 15% .121 23% .148 23% .160 30% .189 28% .239 22% .268 23%
-+DEFACTO[8] .673 2% .715 2% .200 18% .230 20% .220 8% .261 9% .164 4% .200 4% .130 43% .156 41% .277 9% .312 10%
-+AIGC-Synth. .646 2% .685 3% .136 80% .147 95% .170 41% .242 19% .130 21% .165 16% .088 160% .117 125% .234 31% .271 28%
-+MIMLv1 (Conf.) .691 5% .728 4% .302 23% .389 36% .353 48% .422 47% .547 248% .629 228% .397 73% .477 81% .458 50% .529 53%
-+MIMLv2 .684 4% .720 2% .368 50% .443 55% .338 41% .416 45% .562 258% .644 235% .470 105% .548 108% .484 58% .554 60%
-+MIMLv2+O.J .703 7% .742 6% .355 45% .437 53% .358 50% .440 53% .575 266% .656 242% .497 117% .575 119% .498 63% .570 65%
-TruFor B.L. .755 0 .794 0 .363 0 .411 0 .320 0 .377 0 .248 0 .289 0 .296 0 .346 0 .396 0 .443 0
-+PSCC-Synth. .750 1% .798 1% .340 7% .384 7% .325 2% .373 1% .253 2% .298 3% .278 6% .331 5% .389 2% .437 2%
-+DEFACTO[8] .752 0% .791 0% .322 13% .357 15% .214 50% .260 45% .209 19% .247 17% .184 61% .217 59% .336 18% .374 18%
-+AIGC-Synth. .746 1% .791 0% .220 65% .261 57% .213 50% .263 43% .221 12% .267 8% .121 146% .156 122% .304 30% .348 28%
-+MIMLv1 (Conf.) .755 0% .794 0% .385 6% .458 10% .352 9% .408 8% .596 58% .678 57% .340 13% .401 14% .486 18% .548 19%
-+MIMLv2 .767 2% .794 0% .387 6% .458 10% .407 21% .472 20% .618 60% .696 58% .370 20% .422 18% .510 22% .568 22%
-+MIMLv2+O.J .770 2% .811 2% .466 22% .511 20% .439 27% .502 25% .619 60% .700 59% .309 4% .357 3% .521 24% .576 23%
 TABLE VIII: Comparison study of CIML on SPG (left) and SDG (right).
 ’*’ denotes model trained with both SDG and SPG data.
-’Nonzero’ denotes using the non-zero region of the difference map as mask, ’OTSU’ denotes the difference map binarized by OTSU.
+’Nonzero’ denotes using the non-zero region of the difference map as mask.
 ’QES’ denotes the predictions filtered by our QES metric.
@@
 Nonzero .159 .987 .273 .079 DMVN* [40] .553 .357 .434 .276
-OSTU [29] .812 .589 .683 .497 DMVN [26] .438 .568 .495 .317
+OTSU [43] .812 .589 .683 .497 DMVN [40] .438 .568 .495 .317
 DMVN* [40] .329 .618 .430 .276 DMAC* [71] .583 .538 .559 .410
@@
 TABLE IX: Ablation study of CIML on SPG.
-’Forged’ denotes using the forged image as input, ’Auth.’ denotes using the authentic image as input, ’Diff.’ denotes using the absolute difference map between the forged and authentic image as input.
+’Forged’ denotes using the forged image as input, ’Auth.’ denotes using the authentic image as input, ’Diff.’ denotes using the difference map between the forged and authentic image as input.
 Num. Backbone Forged Auth. Diff. P R F1 IoU
@@
 VII-C Experiments on the CAAAv2 Models (CIML task)
-This subsection evaluates the effectiveness of the proposed CAAAv2 in the CIML task by comparing its DASS and Corr-DINO models against other CIML models on SPG and SDG test datasets, respectively.
+This subsection evaluates the effectiveness of the proposed CAAAv2 in the CIML task by comparing its DASS and Corr-DINO models against other CIML models.
+A higher CIML score means more accurate auto-annotation.
 We also conduct detailed ablation studies for in-depth analysis.
@@
 We categorize the forged images in IMD20 into SPG or SDG and randomly split them into training and test sets with a 3:1 ratio.
-CASIAv2 [31] and approximately 1M images synthesized from the COCO [49] dataset are also used for training.
+CASIAv2 [46] and synthetic COCO [67] datasets are also used for training.
 Input images are resized to $\times512$, and a consistent training pipeline is applied across all methods.
@@
 Table VIII presents comparison results for SPG (left) and SDG (right) separately.
-Our method clearly outperforms these previous methods, primarily because our category-aware paradigm significantly reduces task difficulty.
+Our method clearly outperforms these previous methods because our category-aware and prior-feature-denoising paradigm significantly reduces task difficulty.
 As shown in the right side of Table VIII, our Corr-DINO, prior to QES filtering, achieves an IoU of 0.744.
@@
 Notably, DMVN [40] and DMAC [71] models trained on both SPG and SDG data perform worse than those trained exclusively on one category.
-This is attributed to the confusion caused by similar backgrounds in SPG image pairs.
+This is attributed to the confusion caused by similar backgrounds in SPGs, confirming the necessity of our category-aware approach.
 Ablation Study for the DASS Model Input.
-Table IX presents the results, confirming that both the image pair and the image difference map are essential for accurate prediction in SPG.
-The image difference map between forged and authentic images can roughly indicate the forged regions, while the images themselves provide semantic information to help the model denoise the difference map.
-Ablation Study for the Corr-DINO Model Backbone.
-The results are shown in Table X, settings (1-19).
-For the ConvNeXt CNN model, the trainable version performs better.
-For all ViT models, however, frozen versions perform significantly better.
+Table IX presents the results, confirming that both the prior feature difference map and the image pair are essential for accurate prediction in SPG.
+The difference map between forged and authentic images can roughly indicate the forged regions, while the images themselves provide semantic information to help the model denoise the difference map.
+Ablation Study for the Corr-DINO Model Prior Feature Extractor.
+The results are shown in Table X, settings (1-24).
+For ConvNeXt, the trainable version performs better.
+For all ViT-based models, however, frozen versions perform significantly better.
 Since ViT models tend to overfit with limited training data, freezing the backbone alleviates catastrophic forgetting and overfitting.
@@
 The results are shown in the bottom of Table X, setting (25) is the baseline without any proposed modules, setting (29) is the full model with all proposed modules.
-In settings (21, 22, 23), we remove the proposed Multi-Aspect Denoiser, Feature Super Resolution and Learnable Aggregation modules from the full model (24) respectively.
+In settings (26, 27, 28), we remove the proposed Multi-Aspect Denoiser, Feature Super Resolution and Learnable Aggregation modules from the full model (29) respectively.
 Setting (29) significantly outperforms (26).
@@
 This confirms that initial correlation features are too coarse for precise mask predictions, and our Feature Super Resolution module effectively addresses this by reconstructing fine-grained details.
-Finally, Setting (24) is superior to (23) because our learnable aggregation mitigates information loss inherent in rule-based aggregation.
+Finally, Setting (29) is superior to (28) because our learnable aggregation mitigates information loss inherent in previous rule-based aggregation.
 Ablation Study of the QES.
@@
 This confirms that QES effectively filters out unsatisfactory predictions, thereby notably improving annotation quality.
-Table XI confirms QES’s high relevance to the IoU metric without relying on ground truth for evaluation.
+Table XI confirms QES’s high relevance to the IoU metric without relying on ground-truth for evaluation.
+SPG/SDG Classification.
+We evaluate the performance of the SPG/SDG classification model within the proposed CAAAv2 on IMD20 [45] dataset.
+Considering that a very small proportion of SPG image pairs are not spatially aligned, which could negatively impact the prediction’s quality, we also include an additional linear binary classification layer to filter them out.
+The classification results are presented in Table XII.
+It is evident that the voting ensemble of the classification models we used is accurate enough for SPG/SDG classification, as analyzed in Section III.A.
+In practice, to further ensure accuracy, we only keep images with a classification confidence above 99% and discard the rest.
+Our QES also effectively filters the annotations of potentially misclassified samples.
 Our Models are Adequate Auto-Annotators.
 In Table VIII, both our DASS and Corr-DINO achieve high IoU scores.
-Considering the errors in IMD20’s ground truth [54], especially in SPG images, our actual accuracy is higher.
+Considering the errors in IMD20’s ground-truth [74], especially in SPG images, our actual accuracy is higher.
 Clearly, our methods are sufficient for obtaining accurate auto-annotations.
@@
 ‘LA‘ denotes the Learnable Aggregation.
-‘QES‘ denotes filtering the prediction with our QES.
+‘QES‘ denotes filtering the prediction with QES.
 Num. Backbone Freeze MAD FSR LA QES P R F1 IoU
@@
 (9) DINOv1-B [77] L $\times$ .693 .742 .717 .595
-(10) DINOv1-B [57] $\times$ .700 .700 .700 .586
+(10) DINOv1-B [77] $\times$ .700 .700 .700 .556
 (11) DINOv2-S [42] $\times$ $\times$ .739 .653 .693 .584
@@
 (19) DINOv2-L [42] $\times$ .809 .850 .829 .732
+(20) DINOv3-B [78] $\times$ $\times$ .692 .588 .636 .478
+(21) DINOv3-B [78] L $\times$ .690 .753 .720 .598
+(22) DINOv3-B [78] $\times$ .709 .787 .746 .605
+(23) DINOv3-L [78] $\times$ $\times$ .718 .593 .650 .496
+(24) DINOv3-L [78] $\times$ .722 .791 .755 .621
 (25) DINOv2-B [42] $\times$ $\times$ $\times$ $\times$ .644 .605 .624 .469
@@
 IMD20 SDG IoU .744 .825 .842 .865 .889 .912 .924 .957 .967 .968
-TABLE XII: Image manipulation localization performance on SACP [58] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [7] and CASIAv2 [31] datasets for model pretraining
+TABLE XII: Classification results for SDG, SPG and misaligned SPG. ‘Ensemble‘ denotes the ensemble of the models.
+method SDG SPG Misaligned
+P R F P R F P R F
+DiNAT [79] .992 .992 .992 1 .996 .998 .867 .929 .897
+SwinTrans [80] .996 .981 .988 1 .996 .998 .737 1 .849
+ConvNeXt [68] .996 .992 .994 1 .996 .998 .875 1 .933
+Ensemble .996 1 .998 1 .996 .998 1 1 1
+TABLE XIII: Image manipulation localization performance on SACP [81] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [19] and CASIAv2 [46] datasets for model pretraining
 Num. Methods Pretraining data IoU F1
-(1) DFCN [59] Image-Net [60] .466 .607
-(2) MVSS-Net [20] Image-Net [60] .401 .534
-(3) PSCC-Net [37] Image-Net [60] .482 .620
-(4) RRU-Net [43] Image-Net [60] .517 .651
-(5) CFL-Net [61] Image-Net [60] .433 .571
-(6) DTD [62] Image-Net [60] .588 .712
-(7) TIFDM [63] Image-Net [60] .576 .703
-(8) Web-IML (Ours) Image-Net [60] .773 .862
+(1) DFCN [82] ImageNet [83] .466 .607
+(2) MVSS-Net [35] ImageNet [83] .401 .534
+(3) PSCC-Net [56] ImageNet [83] .482 .620
+(4) RRU-Net [60] ImageNet [83] .517 .651
+(5) CFL-Net [84] ImageNet [83] .433 .571
+(6) DTD [85] ImageNet [83] .588 .712
+(7) TIFDM [86] ImageNet [83] .576 .703
+(8) Web-IML (Ours) ImageNet [83] .773 .862
 (9) Web-IML (Ours) ADE20k [87] .774 .863
@@
 (11) Web-IML (Ours) (10) + DEFACTO [20] .767 .856
-(12) Web-IML (Ours) (10) + PSCC Synthetic [37] .762 .853
+(12) Web-IML (Ours) (10) + PSCC-Synthetic [56] .762 .853
 (13) Web-IML (Ours) (10) + AIGC Synthetic [21] .756 .851
@@
 This subsection evaluates the generalization of the Web-IML model and the universal applicability of our web-supervision method on an important downstream task: document IML.
-We fine-tune our Web-IML under various pretraining settings on the handcrafted document IML benchmark SACP [58], using the same training and test configurations as prior works [63].
-Our Web-IML notably outperforms prior methods.
-Moreover, Table XII shows that settings (11, 12, 13), which incorporate additional training data from DEFACTO [8], PSCC [37], and OpenSDID [48] into the baseline (Setting 10), perform worse than the baseline.
-This confirms that simply increasing the scale and diversity of pretraining data does not lead to better generalization.
-In contrast, the model pretrained on our web supervision data (15) shows a significant 2.2 point IoU improvement over the baseline (10), and is 0.8 points higher than the model pretrained on MIMLv1 (14) from our conference paper.
-These results demonstrate that the cheap data from our web supervision method is of sufficient quality to enable models to learn semantic-agnostic forensic features and generalize across unseen real-world IML tasks.
-VIII Conclusion
-We propose CAAAv2, a novel category-aware framework that solves the constrained image manipulation localization problem by separately handling Shared Donor Group (SDG) and Shared Probe Group (SPG) image pairs through Difference-Aware Semantic Segmentation and Correlation DINO with frozen DINOv2 backbone, dramatically reducing overfitting while enabling high-quality automatic annotation.
-With CAAAv2 and a novel metric QES, we auto-annotate manually forged images from the web and harvest a large-scale, diverse, high-quality dataset MIMLv2, comprising 246,212 manually forged images with pixel-level annotations.
-This dataset considerably addresses the data scarcity issue in IML.
-Furthermore, we propose Object Jitter, an effective data generation method that leverages authentic web images to improve model performance in challenging cases.
-To better leverage web-scale supervision, we also introduce Web-IML, a new model for IML.
-Experiments confirm the strong generalization ability of our method, which achieves unprecedented performance gains and surpasses previous state-of-the-art methods by significant margins.
-We believe this work establishes a scalable framework for continuously expanding training data as new manually forged images emerge online.
-Its universal applicability is demonstrated across multiple benchmarks.
-This research transforms image forensics, shifting reliance from limited, expensive handcrafted datasets to abundant, continuously growing web resources, thus paving the way for more robust and generalizable IML systems.
-Figure 12: Qualitative comparison on image manipulation localization.
+We fine-tuned the Web-IML model, pretrained under various configurations, on the handcrafted real-world document IML benchmarks, SACP [81] and RTM [88], adopting the same training and test configurations as previous studies [86, 88].
+Across both benchmarks, our Web-IML model consistently and notably outperforms existing methods.
+A key observation from Tables XIII and XIV is that settings 11, 12, and 13, which augmented the baseline (Setting 10) with additional pretraining data from DEFACTO [20], PSCC [56], and OpenSDID [21], respectively, exhibited degraded performance.
+This confirms that merely increasing the scale and diversity of pretraining data does not lead to improved real-world generalization.
+In contrast, the model pretrained on our web-supervision data (Setting 15) demonstrates significant improvement.
+It achieved 2.2 points higher IoU on SACP and 4.7 points on RTM compared to the baseline (Setting 10).
+Furthermore, it also surpassed the model pretrained on MIMLv1 (Setting 14) by 0.8 points higher IoU on SACP and 1.5 points on RTM.
+These results demonstrate that the cost-effective data created by our web-supervision method possesses sufficient quality and relevance to enable models to learn robust semantic-agnostic forensic features, thereby generalizing effectively across diverse and unseen real-world IML tasks.
+TABLE XIV: Image manipulation localization performance on RTM [88] dataset. ‘Forgery Baseline‘ denotes using the tampCOCO [19] and CASIAv2 [46] datasets for model pretraining.
+Num. Methods Pretraining data IoU F1
+(1) RRU-Net [60] ImageNet [83] .037 .072
+(2) PSCC-Net [56] ImageNet [83] .033 .064
+(3) MVSS-Net [35] ImageNet [83] .051 .097
+(4) CAT-Net [19] ImageNet [83] .113 .203
+(5) DTD [85] ImageNet [83] .065 .122
+(6) Mask2Former [89] ADE20k [87] .124 .172
+(7) ASCFormer [88] ADE20k [87] .198 .329
+(8) Web-IML (Ours) ImageNet [83] .192 .323
+(9) Web-IML (Ours) ADE20k [87] .207 .343
+(10) Web-IML (Ours) Forgery Baseline .194 .325
+(11) Web-IML (Ours) (10) + DEFACTO [20] .186 .313
+(12) Web-IML (Ours) (10) + PSCC-Synthetic [56] .189 .318
+(13) Web-IML (Ours) (10) + AIGC Synthetic [21] .185 .311
+(14) Web-IML (Ours) (10) + MIMLv1 [24] .226 .369
+(15) Web-IML (Ours) (10) + Web Supervision (Ours) .241 .389
+Figure 10: Qualitative evaluation on image manipulation localization.
+From left to right: input image, Ground-Truth (abbreviated as GT), RRU-Net (RRU), PSCC-Net (PSCC), MVSS-Net (MVSS), IF-OSN (OSN), CAT-Net (CAT), TruFor (T.F.), SparseViT (S.ViT), APSC-Net (APSC) and ours.
+Figure 11: Both Web-IML and previous methods sometimes struggle to produce precise mask when the forged region exhibits few visual artifacts.
+From left to right: input image, Ground-Truth (abbreviated as GT), RRU-Net (RRU), MVSS-Net (MVSS), IF-OSN (OSN), CAT-Net (CAT), TruFor (T.F.), SparseViT (S.ViT), APSC-Net (APSC) and ours.
+VIII Limitations
+Although our QES metric is highly effective in ensuring overall annotation quality, a small number of correct annotations may also be filtered out.
+In addition, our Web-IML is designed to maximize accuracy by leveraging web-scale supervision with a medium-sized model (140M).
+As a result, it is not highly computationally efficient (achieving 10 frames per second on a single 3090 GPU).
+Figure 11 presents some failure cases, showing that both Web-IML and previous methods may struggle to produce precise masks when the manipulated region exhibits very few visual artifacts.
+IX Conclusion
+In this paper, we addressed the critical challenge of data scarcity in image manipulation localization through a scalable web-supervision approach.
+At the core of our approach is CAAAv2, a novel category-aware and prior-denoising auto-annotation paradigm.
+This innovative design strategically separates manipulations into Shared Donor Group and Shared Probe Group, effectively converting the challenging forensic task into a prior-denoising task.
+This dramatically reduces task complexity and mitigates overfitting.
+Leveraging this automatic annotation engine and our proposed QES metric for quality control, we constructed MIMLv2, a diverse, high-quality dataset of 246,212 manually forged images—over 120× larger than existing handcrafted datasets.
+We further enhanced this web-scale supervision with Object Jitter, a novel technique for generating challenging training examples from authentic images.
+To fully capitalize on this vast data, we designed Web-IML, a new model featuring multi-scale perception and self-rectification modules.
+Extensive experiments demonstrate that our web-supervision method provides unprecedented performance gains across multiple models on a wide array of real-world forgery benchmarks.
+Additionally, Web-IML establishes a new state-of-the-art, significantly outperforming previous methods and demonstrating remarkable generalization.
+We believe that this work establishes a scalable framework for continuously expanding training data as new manually forged images emerge online.
+This research transforms image forensics, shifting reliance from limited, expensive handcrafted datasets to abundant, continuously growing web resources, thus paving the way for more robust image manipulation localization systems.